### v1

# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

Abstract
Owing to its outstanding parallel computing capabilities, quantum computing (QC) has been a subject of continuous attention. With the gradual maturation of QC platforms, it has increasingly played a significant role in various fields such as transportation, pharmaceuticals, and industrial manufacturing, achieving unprecedented milestones. In modern society, wireless communication stands as an indispensable infrastructure, with its essence lying in optimization. Although artificial intelligence (AI) algorithms such as Reinforcement Learning (RL) and mathematical optimization have greatly enhanced the performance of wireless communication, the rapid attainment of optimal solutions for wireless communication problems remains an unresolved challenge. QC, however, presents a new alternative. This paper aims to elucidate the fundamentals of QC and explore its applications in wireless communications and networking. First, we will provide a tutorial on QC, covering its basics, characteristics, and popular QC algorithms. Next, we will introduce the applications of QC in communication and networking, followed by its applications in miscellaneous areas such as security and privacy, localization and tracking, and video streaming. Finally, we will discuss remaining open issues before concluding.
Index Terms:
quantum computing, quantum algorithm, wireless communications and networking, resource allocation, scheduling, offloading, edge computing, security and privacy, localization, video streaming, reinforcement learning
I Introduction
I-A Background and Motivation
Quantum computing (QC) [1, 2], a revolutionary approach to processing information based on quantum mechanics principles, shows great potential for transforming various industries. Unlike traditional computers that use bits in a binary state (0 or 1), quantum computers apply qubits that can exist in multiple states simultaneously, allowing for much more powerful computation. This technology is expected to greatly improve cryptography by breaking current encryption methods and enabling the development of more secure ones. Additionally, QC is highly effective in solving optimization problems, which can benefit applications like route planning and managing supply chains. It also has applications in fields such as drug discovery, materials science, and simulating quantum systems, offering faster insights into complex interactions at the molecular level. As of December 2023, the investment for quantum market reaches to \$6.7B with 261 start-ups, and potential quantum technology market size by 2040 is estimated by \$173B [3], fueled by substantial investments from governments and industries harnessing the transformative potential of this cutting-edge technology.
On the other hand, wireless communications and networking play a foundational role in modern connectivity, facilitating the smooth exchange of data across devices and supporting a wide range of applications, from mobile networks and WiFi to IoT devices and satellite communications [4]. Optimization is crucial in wireless networks, especially as we move towards the era of 6G [5]. With the increasing demand for higher data rates, lower latency, and greater device density, optimizing spectrum usage, reducing interference, and improving signal processing algorithms become even more critical. 6G, the next frontier in wireless technology, is poised to offer unprecedented speeds, ultra-low latency, and innovative applications [6]. Effectively optimizing these advanced wireless networks will be essential for fully realizing the potential of 6G, ensuring efficient and reliable communication in an increasingly interconnected world.
Additionally, the interaction among the aforementioned enabling technologies holds the potential for significant performance enhancements, encompassing increased capacity, enhanced energy efficiency, and improved reliability. For instance, incorporating edge intelligence for micro-service orchestration, as demonstrated in [7], can bolster network throughput and reduce service delay in next-generation networks. Leveraging QC algorithms for optimization tasks can accelerate this process. As wireless networks expand to encompass intelligent devices, such as autonomous vehicles, UAVs and satellites, the complexity of problems increase accordingly [8]. In particular, optimization tasks in wireless communications and networking, like channel assignment, power allocation, edge computing, and video streaming demand substantial computational resources. While conventional solutions often rely on local servers, cloud infrastructure, or GPU servers, they still struggle to meet the growing research demands. QC techniques offer a promising approach to addressing these challenges by leveraging its ability to handle complex optimization scenarios efficiently, such as slicing [9] and Ultra-Reliable Low Latency Communications (URLLC) [10] in 6G.
Given this landscape, our survey aims to explore the intersection of QC and wireless communications and networking, providing insights into the latest developments, challenges, and opportunities in this rapidly evolving field. We explore the application of QC techniques to address optimization challenges in wireless networks, offering a comprehensive overview of current research trends and future directions. Through this survey, we seek to shed light on the transformative potential of QC in reshaping the landscape of wireless communications and networking.
I-B Related Surveys
While previous surveys, primarily concentrated on quantum communications [11, 12, 13], have explored the applications and advantages of integrating QC into wireless networks from various applications [14, 15, 16, 17], a comprehensive survey on QC-enabled communication and networking is still lacking. The first survey [14] investigated the utilization of QC in addressing challenges within wireless communication systems. It provided an overview of QC fundamentals using linear algebra, showing its potential for enhancing wireless communication systems as proposed in existing literature. The survey extensively explored various optimization problems encountered in both the physical and network layers of wireless communications, offering comparisons between classical and quantum-assisted solutions. However, the survey from a few years back lacked coverage of AI-powered QC algorithms such as Quantum Neural Networks (QNNs) and Quantum Deep Reinforcement Learning (QDRL). Moreover, it did not address particular concerns within wireless networks like edge computing and video streaming, which are especially relevant in the context of 5G/6G networks.
TABLE I: Summary of related surveys on applications of QC in wireless networks
Tutorial on QC QC in Communication QC in Networking Others
Refs Year BQ CQ AQ PA CA UA R TO CC SP LT VS
[14] 2018 PD
[15] 2019 PD PD PD PD PD PD PD PD
[11] 2021
[16] 2022 PD PD PD PD PD
[17] 2023 PD PD
[12] 2023
[13] 2023 PD
[18] 2024 PD PD PD PD PD
ours 2024
BQ: Basic of QC; CQ: Conventional QC; AQ: AI-driven QC; PA: Power Allocation; CA: Channel Assignment; UA: User Association; R: Routing; TO: Task Offloading; CC: Content Caching; SP: Security and Privacy; LT: Localization and Tracking; VS: Video Streaming; PD: Partially Addressed.
AI-driven QC algorithms have been explored in several studies, including [15, 16, 17, 18]. However, in [15], the discussion on conventional and AI-driven QC algorithms remained largely conceptual, with limited elaboration on specific algorithms like QDRL. Topics such as task offloading and content caching in wireless networks were also presented conceptually. Another brief survey on QC algorithms for 6G networks, focusing on quantum-inspired machine learning applications, can be found in [16]. This survey provided a brief overview of both conventional and AI-driven algorithms, including the Shor algorithm, Grover algorithm, QRL, and QNNs. It briefly discussed a few topics mainly related to security and resource allocation for 6G networks. The fundamental principles of QC, as well as conventional and AI-driven QC algorithms, were relatively described in [17]. However, the paper provided only a limited introduction to their applications in wireless networks, focusing mainly on power allocation and channel assignment. In [18], the applications of QC were reviewed in details, with a focus on quantum cloud computing. The paper discussed the latest advancements in quantum cloud technology, covering various cloud-based models, platforms, and recently developed technologies and software use cases. It also examined several applications of quantum cloud computing, such as resource management, quantum serverless computing, and addressing security and privacy concerns. Table I presents a comparative summary of the referenced surveys, underscoring that the most recent ones fail to provide a comprehensive overview of the QC applications in wireless networks. This observation highlights a gap in the existing literature.
I-C Contributions
We take a broad approach aimed at a wide audience to show the QC advantages, offering the following contributions.
  • Our survey offers a comprehensive introduction to fundamental quantum mechanics concepts, such as quantum measurement, essential for developing a deep understanding of QC.
  • We thoroughly explain all aspects of QC, ensuring that even readers without a background in quantum mechanics can easily grasp the foundational principles of this emerging technology.
  • We encompass all QC algorithms, ranging from classical approaches like Quantum Approximate Optimization Algorithm (QAOA), to quantum-inspired metaheuristic methods like Quantum Particle Swarm Optimization (QPSO), and even to AI-driven techniques such as QDRL.
  • We examine recent advancements in the application of QC algorithms to address various communication and networking challenges in wireless networks, spanning power allocation, channel assignment, user association, routing, edge computing/task offloading, content caching, security and privacy, localization and tracking, video streaming, and more. Additionally, we identify key challenges and open problems for future research to fully utilize the potential of QC.
I-D Literature Categorization and Paper Organization
Figure 1: Paper organization
We have adopted the following methodology to ensure a comprehensive and in-depth review of applications of QC algorithms. We primarily organize the papers from two dimensions: the research issues and category of QC algorithms. We divide the research issues in wireless networks into two parts: communications from the perspective of the physical layer, where issues typically arise within a single hop, and networking from the perspective of the routing layer, where issues typically arise within multiple hops. The former mainly considers topics including channel assignment, power allocation, and user association, while the latter mainly considers routing, task offloading, content caching, and among others. For each issue, we consider its network characteristics such as the presence of multiple cells in communications and mobility in networking. Additionally, we divide QC algorithms into two categories: conventional QC algorithms, such as classical QAOA and quantum-inspired heuristics like QPSO and Quantum Genetic Algorithm (QGA), and AI-driven algorithms, such as QNNs and QDRL. The paper organization is described in Fig. 1.
TABLE II: Acronyms and Abbreviations
Acronym Full form
QC Quantum Computing
AI Artificial Intelligence
QAOA Quantum Approximate Optimization Algorithm
D2D Device to Device
BS Base Station
NISQ Noisy Intermediate-Scale Quantum
QA Quantum Annealing
QPSO Quantum Particle Swarm Optimization
QML Quantum Machine Learning
QUBO Quadratic Unconstrained Binary Optimization
MDP Markov Decision Process
QRL Quantum Reinforcement Learning
QDRL Quantum Deep Reinforcement Learning
QGA Quantum Genetic Algorithm
QNN Quantum Neural Network
SVM Support Vector Machine
PQC parameterized Quantum Circuits
QGLM Quantum Generative Learning Model
NOMA Non-orthogonal Channel Multiple Access
MIMO Multiple Input Multiple Output Antennas
FDMA Frequency-division Multiple Access
OFDMA Orthogonal Frequency-division Multiple Access
RSU Road-side Unit
IRS Intelligent Reflecting Surface
MEC Mobile Edge Computing
VQE Variational Quantum Eigensolver
QMARL Quantum Multiple Agent Reinforcement Learning
ILP Integer Linear Programming
QEA Quantum Evolution Algorithm
URLLC Ultra-Reliable Low Latency Communications
II A Tutorial of QC
II-A The Postulates of Quantum Mechanics
QC is grounded in the principles of quantum mechanics and unequivocally adheres to its fundamental rules. In this section, we will outline the essential principles of quantum mechanics [19, 20], providing the requisite foundation for the subsequent discussion. This explanation is crucial in establishing a mathematical framework for comprehending QC, thereby facilitating a deeper understanding of this intricate field for readers without a background in physics.
Postulate 1: Every enclosed physical system possesses a corresponding Hilbert space, wherein the system state can be precisely depicted by a unit vector.
The Hilbert space, as delineated in the given postulate, corresponds to a complete complex linear space, characterized by a defined inner product, pertinent to QC since it does not involve infinite spatial dimensions. A collection of $n$ linearly independent vectors is capable of spanning an n-dimensional linear space. When these vectors are both normalized and mutually orthogonal, they constitute what is known as orthonormal bases for the aforementioned space. A unique unit vector, designated as a ’state’, encapsulates the description of a specific physical system. To articulate this with greater precision and mathematical rigor, the state of a physical system is depicted by a unit vector. This unit vector can be linearly decomposed using any complete set of orthonormal bases within the linear space, offering a mathematically precise representation of the physical system in question.
Physicist Dirac creates a set of symbols to describe system states [21]. The symbol $\ket{.}$ is called ket, corresponding to a column unit vector in the space. Thus, a specific system state, denoted by a quantum state |$\ket{\phi}$ , can be represented linearly by a set of orthonormal bases {|$\{\ket{a_{1}},\ket{a_{2}},\dots,\ket{a_{n}}\}$.
|$\ket{\phi}=\sum_{i=1}^{n}c_{i}\ket{a_{i}},\quad st.\sum_{i}|c_{i}|^{2}=1.$ (1)
The algebra reveals that the orthonormal bases of a Hilbert space are not unique; thereby, a quantum state can be described by different bases. This property is known as representation transformation in quantum mechanics and is essentially equivalent to the basis transformation.
The conjugate transpose of |$\ket{\phi}$ is written as $\bra{\phi}$ |, corresponding to a row vector in the Hilbert space. The inner product of two vectors is defined by the dot product of the row vector $\bra{\psi}$ | and the column vector |$\ket{\phi}$, represented as $\bra{\psi}\ket{\phi}$.
Postulate 2: The evolution of a closed quantum system |$\ket{\phi}$ can be described by the Schrodinger equation:
$i\hbar\frac{d\ket{\phi(t)}}{dt}=H\ket{\phi(t)},$ (2)
where $\hbar$ is reduced Plank constant, $H$ is the Hamiltonian of the system, mathematically expressed as a Hermite operator, i.e., $H^{\dagger}=H$.
In general, finding the Hamiltonian to characterize a particular physical system is difficult (requiring substantial experimental results), but fortunately this is not the primary concern for the involved theoretical framework. In the discussions of QC and quantum information, we always assume that $H$ is known. Thus, if the system in the state of time $t_{0}$ is |$\ket{\phi(t_{0})}$, then the system in the $t_{1}$ moment can be solved by the Schrodinger equation in eq. (3),
|$\ket{\phi(t_{1})}=exp\left[\frac{-iH(t_{1}-t_{0})}{\hbar}\right]\ket{\phi(t_{0})}=U(t)\ket{\phi(t_{0})},$ (3)
where $U(t)$ is the evolutive operator from $t_{0}$ to $t_{1}$, defined as:
$U(t)=exp\left[\frac{-iH(t_{1}-t_{0})}{\hbar}\right].$ (4)
It is easy to verify that $U(t)$ is a unitary operator, i.e., $U(t)^{\dagger}U(t)=I$. In many cases, for the convenience of discussion, eq. 4 can be simplified in eq. (5),
|$\ket{\phi_{1}}=U\ket{\phi_{0}}.$ (5)
Postulate 3: A general quantum measurement is described by a family of linear operators $\{M_{m}\}$. These operators act on the state space of the system to be measured, satisfying the completeness condition $\sum_{m}M_{m}^{\dagger}M_{m}=I$, and the indicator $m$ represents the possible measurement results in the experiment.
Assuming that the state of the system before measurement is |$\ket{\phi}$, the probability of observing the result $m$ is in eq. (6),
$p(m)=\bra{\phi}M_{m}^{\dagger}M_{m}\ket{\phi}.$ (6)
The post-measurement state is in eq. (7),
[…]
Postulate 4: The state space of a composite physical system is the tensor product of the state space of a sub-physical system.
If the index number of subsystems is $1$ to $n$, and the state of subsystem-$i$ can be represented as |$\ket{\phi_{i}}$, then the state of the entire system is |$\ket{\phi_{1}}\otimes\cdots\otimes\ket{\phi_{n}}$. The notation $\otimes$ represents the tensor product, which is a way of combining smaller vector spaces together to form a larger vector space. Its operation rules can be denoted by Kronecker multiplication in eq. (8),
[…]
where A and B are $m\times n$ and $p\times q$ matrix, respectively.
II-B The Characteristics of QC
1) Properties of Qubits
In the domain of classical computing, the bit, with possible states 00 or $1$, serves as the fundamental unit of information. This binary state system can represent a physical system with two clearly distinguishable states, exemplified by the widely utilized high and low-level electrical signals.
Conversely, in QC, the basic information unit is the quantum bit (qubit). A qubit has the capability to physically manifest in a variety of forms. Examples of this include the ground state and the first excited state of an electron within a hydrogen atom; the $+\frac{1}{2}$ and $-\frac{1}{2}$ components of a proton’s spin in any given direction; or the clockwise and anti-clockwise circular polarization of light. A notable distinction of qubits, as opposed to classical bits, is their ability to exist not merely in a binary state of ‘dimorphism’ but also in any intermediary phase between these two states, a phenomenon known as linear superposition. Adhering to Postulate 1, a qubit can be mathematically represented as a unit vector within a two-dimensional Hilbert space, as delineated in eq. (9)
|$\ket{\phi}=\alpha\ket{0}+\beta\ket{1}\quad st.\ |\alpha|^{2}+|\beta|^{2}=1,$ (9)
where the complex numbers $\alpha$ and $\beta$ satisfy the normalization condition, and |$\ket{0}$ and |$\ket{1}$ are orthonormal basis $\begin{pmatrix}1\\ 0\end{pmatrix}$ and $\begin{pmatrix}0\\ 1\end{pmatrix}$. Thereby, |$\ket{\phi}$ is represented as $\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$.
There are several distinctive properties of qubits:
• The amount of represented information increases exponentially with the number of qubits.
A qubit is a unit vector in a two-dimensional Hilbert space. The representation of two qubits can be derived by (Kronecker product eq. 8) of two single qubits. For example, |$\ket{\phi_{1}}=\alpha_{1}\ket{0}+\beta_{1}\ket{1}$, |$\ket{\phi_{2}}=\alpha_{2}\ket{0}+\beta_{2}\ket{1}$, then two qubits:
[…]
where |$\ket{00}=\begin{pmatrix}1\\ 0\\ 0\\ 0\end{pmatrix}$, |$\ket{01}=\begin{pmatrix}0\\ 1\\ 0\\ 0\end{pmatrix}$, |$\ket{10}=\begin{pmatrix}0\\ 0\\ 1\\ 0\end{pmatrix}$, and |$\ket{11}=\begin{pmatrix}0\\ 0\\ 0\\ 1\end{pmatrix}$, form a set of orthonormal bases for a four-dimensional Hilbert space. Obviously, $|\alpha_{1}\alpha_{2}|^{2}+|\alpha_{1}\beta_{2}|^{2}+|\beta_{1}\alpha_{2}|^{2}+|\beta_{1}\beta_{2}|^{2}=1$, and |$\ket{\phi_{1}\phi_{2}}$ is a normalized vector in a four-dimensional Hilbert space. By analogy, the extension of $n$ qubits |$\ket{\phi_{1}\phi_{2}\cdots\phi_{n}}$ can be expressed in eq. (11),
|$\ket{\phi_{1}\phi_{2}\cdots\phi_{n}}=\sum_{i=0}^{2^{n}-1}c_{i}\ket{i}\qquad st.\ \sum|c_{i}|^{2}=1,$ (11)
Suppose a quantum register in a quantum computer stores $n$ qubits. According to eq. (10), the quantum register stores every number from 00 to $12^{n}-$ 1 simultaneously with the probability of $|c_{i}|^{2}$. In other words, a $n$-qubit quantum register holds $2^{n}$ binary numbers.
• Probability statistical properties of measurement
While $n$ superposed qubits can concurrently represent $2^{n}$ binary numbers, there is generally no methodology to simultaneously extract each binary number concealed within the superposition state. More specifically, one cannot accurately determine the probability amplitude $c_{i}$ (eq. 11) of each individual binary number for an unknown $n$-qubit. Our primary tool for understanding an unknown quantum state is through measurement or observation. However, each observation only yields a fragment of information about the initial state since the state collapses upon measurement, leading to significant data loss.
To illustrate, consider a beam of light emitting a mixture of horizontally and vertically polarized photons |$\ket{\phi}=\alpha\ket{\uparrow}+\beta\ket{\rightarrow}$. Once this beam passes through a horizontal polarization filter, only vertically polarized photons $\ket{\uparrow}$ can be observed. This observation results in |ϕ|ketitalic-$\ket{\phi}\Rightarrow\ket{\uparrow}$, with the frequency of observing vertically polarized light (or vertically polarized photons) approximating $|\alpha|^{2}$. Similarly, if this light is subjected to a vertical polarization filter, the observable light will be horizontally polarized with a probability of $|\beta|^{2}$.
The methodology of this aforementioned observation characterizes the process of projective measurement, a specific case of the more general measurement process described by Postulate 3. This type of measurement, alternatively known as Von Neumann measurement, is extensively employed across various studies. Regardless of physical circumstances, this can be abstracted as an observable or a Hermitian operator $M$, presented in the spectral decomposition form as shown below.
$M=\sum_{m}mP_{m},$ (12)
where the projective operator $P_{m}$ is the projection onto the eigenspace $M_{m}$ corresponding to the eigenvalue $m$, $P_{m}P_{m^{\prime}}=\delta_{mm^{\prime}}P_{m}$, and satisfies completeness condition $\sum_{m}P_{m}=I$.
The measurement outputs are eigenvalues $m$. When measuring the state |$\ket{\phi}$, the probability of getting a certain $m$ is in eq. (13),
$p(m)=\bra{\phi}P_{m}\ket{\phi}.$ (13)
After the measurement, |$\ket{\phi}$ collapses to eq. (14),
$\frac{P_{m}\ket{\phi}}{\sqrt{p(m)}},$ (14)
where if there is no ambiguity in the context, it is said that the measurement is performed in basis |$\ket{m}$; thereby, the projective operator $P_{m}$ can be written as $P_{m}=\ket{m}\bra{m}$ |.
From the preceding exposition, it is evident that generally, post-measurement qubits |$\ket{\phi}$ alter their state to a particular basis |$\ket{m}$ with probability $p(m)$, unless the basis of measurement coincides with |$\ket{\phi}$ and |ϕketsuperscriptitalic-$\ket{\phi^{\bot}}$. This characteristic is in stark contrast to classical bits, which, irrespective of the form of measurement (observation) or the frequency with which measurements are performed, remain in their initial state. For a state such as |$\ket{\phi}=\sum_{i=0}^{2^{n}-1}c_{i}\ket{i}$, executing a projective measurement $P_{i}=\ket{i}\bra{i}$ | will result in |$\ket{i}$ with a probability of $|c_{i}|^{2}$.
• Quantum states cannot be cloned
A notable attribute of quantum states is that only partial information can be gleaned through a single measurement. It is impossible to ascertain the complete information of a quantum state by repetitively cloning and measuring it. A simplified proof supporting this conclusion is provided herein, with additional specifics available in [22].
Assume a unitary operation $U$ is able to replicate quantum state |$\ket{\phi}$ and substitute quantum resource |$\ket{s}$ with |$\ket{\phi}$:
[…]
Accordingly, for |$\ket{\phi}$ and |$\ket{\psi}$:
[…]
Apply the inner product on both sides of the above two equations, we obtain:
[…]
The resulting equation suggests $x=x^{2}$, whereby $x=0$ or $x=1$ ($\bra{\phi}\ket{\psi}=x$). This necessitates that either |$\ket{\phi}=\ket{\psi}$ or |$\ket{\phi}$ and |$\ket{\psi}$ are orthogonal. As a consequence, a quantum cloning device can only clone mutually orthogonal states, thereby proving that a universal quantum cloning machine cannot exist. In essence, a unitary operation that flawlessly duplicates an unknown quantum state does not exist unless the qubit has degenerated into a classical bit.
2) Quantum Parallelism
The most compelling characteristic of QC is its parallel computing capacity, which arises from the superposition and entanglement of quantum states. As per Postulate 2, QC fundamentally revolves around the unitary evolution of quantum states. Mathematically, unitary evolution can be depicted by a unitary matrix $U_{f}$, which adheres to eq. (18),
$U_{f}U_{f}^{\dagger}=I,$ (18)
where $U_{f}^{\dagger}$ signifies the conjugate transpose of $U_{f}$ and $I$ the identity matrix.
Consider a classical computational function $f:x\rightarrow f(x)$. This function $f$ can be transformed into an associated QC function $U_{f}$ in terms of polynomial cost [23]. Thus, the newly converted $U_{f}$ will implement $U_{f}:\ket{x}\rightarrow\ket{f(x)}$. Suppose an $n$-qubit register stores qubits with initial values |$\ket{00\cdots 0}$. Firstly, $H^{\otimes n}$ is applied in eq. (19) to the qubits, consequently generating a superposition state comprising all binary numbers between 00 and $12^{n}-$ 1 with an equal probability of $\frac{1}{2^{n}}$.
$H^{\otimes n}\ket{00\cdots 0}=\frac{1}{\sqrt{2^{n}}}\sum_{x=0}^{2^{n}-1}\ket{x}.$ (19)
Subsequently, $f(x)$ is computed by $U_{f}$ as per eq. 20,
$U_{f}\!\!\left(\!\!\frac{1}{\sqrt{2^{n}}}\!\sum_{x=0}^{2^{n}-1}\ket{x}\!\!\right)\!\!=\!\!\frac{1}{\sqrt{2^{n}}}\sum_{x=0}^{2^{n}-1}U_{f}\ket{x}\!\!=\!\!\frac{1}{\sqrt{2^{n}}}\!\sum_{x=0}^{2^{n}\!-\!1}\ket{f(x)}$ (20)
[…]
This feature of QC, often termed quantum parallelism, is the fundamental reason for its superiority over classical computation. Despite quantum parallel operations typically generating a simultaneous output of $2^{n}$ results, it remains a challenge to read all |$\ket{f(x)}$ states in one go. The sole approach to extracting information from quantum states is through measurement, but each measurement only retrieves a fraction of the resultant information and triggers the collapse of the quantum state. To amass all the information pertaining to the final result, repeated measurements at varying bases (typically exponential) are required, thereby necessitating considerable resources.
II-C The Description of Quantum Algorithm
1) Overview of quantum algorithm capabilities
The preceding exposition of quantum parallelism results in distinctly opposing observations. On one hand, quantum parallelism holds the potential to tremendously boost computing efficiency. Conversely, it can be costly to acquire comprehensive and accurate solutions. Thus, it prompts the question, do quantum algorithms truly offer an advantage?
To answer this question, we should investigate quantum algorithms from the perspective of computational theory through more fundamental problems. How can it be ascertained whether a problem can be addressed using a quantum algorithm? How should a quantum algorithm be described? What’s the mechanism for gauging the computational efficiency and the resource utilization of a quantum algorithm?
The exploration of these questions originated with Benioff’s work [24], wherein he initially coined the quantum Turing machine model. This model was later elaborated by Deutsch [25, 26] and Yao [27], who introduced the quantum circuit model, proving its equivalent standing with the Turing model. Subsequent studies by Bernstein [28] and Knill [29] formed the foundation of quantum computational complexity theory, developing links with classical computational complexity. These classic documents provide the answer:
• A problem unsolvable by classical computation also remains unsolvable by quantum algorithms.
The Church-Turing proposition claims that any deterministic computational process, regardless of its complexity, can be decided by a Turing machine. From an abacus to a supercomputer, any instrument capable of computation can be simulated by a Turing machine. Hence, all computations can be equated to effective simulations on a Turing machine. To determine whether a mathematical problem is solvable depends on two conditions: whether the input of the problem is countable and whether the Turing machine would stop. Besides the well-known ‘halting problem over Turing machines’, several problems remain unsolvable [30], such as ascertaining if two topological spaces are topologically equivalent (homeomorphic). Can these ostensibly solvable, yet inherently unsolvable problems be resolved on a quantum computer? The answer is negative. A quantum Turing machine is incapable of resolving an issue that a Turing machine cannot, implying that from the standpoint of computability, both the quantum Turing machine and the Turing machine have the same ability. Consequently, problems that are unsolvable through classical methods also remain irresolvable by quantum algorithms.
• Quantum algorithms can solve problems that cannot be solved efficiently in classical computation.
The ability to solve problems and to solve problems efficiently, are distinct notions. Problems that can be efficiently solved refer to issues resolvable through an algorithm that incurs a polynomial cost relative to the problem scale. It is generally agreed that quantum algorithms can efficiently solve some problems that classical computation cannot address efficiently. While this assertion has not been distinctly proven, substantial evidence supports its validity. Shor’s algorithm [31] is a prime example, efficiently solving large integer factorization and discrete logarithm in $\mathcal{O}(poly(n))$, while no efficient classical algorithms exist to date. Quantum Turing machines deal with vectors, whereas Turing machines handle character strings. Additionally, the transition relations are expected to be more plentiful given the larger number of legal quantum gates compared to logic gates. Accordingly, it is widely accepted that the relation $P\subseteq BQP$ is valid, where BQP represents the class of problems that quantum algorithms can solve efficiently.
• Quantum circuits serve as an efficient tool for the depiction and analysis of quantum algorithms.
The Quantum Turing machine is generally cumbersome for describing quantum algorithms and is more suitable for analyzing and positing the inherent relationship between quantum and classical algorithms. Similar to steering clear of high-level applications in assembly language, a quantum circuit stands as a practical tool for the analysis and description of a quantum algorithm. Yao’s work attests to the equivalence of quantum circuits and quantum Turing machines [27], while Barenco et al. furnished a basic method of analyzing quantum circuits and unveiling the elementary gates within them [32]. These elementary gates form universally complete subsets for describing quantum algorithms. In simple terms, a corresponding quantum circuit can always be constructed to analyze and measure the efficiency and costs of even the most complex quantum algorithms by evaluating circuit depth and gate counts.
Besides quantum circuits, two higher-level ways exist for the description of quantum algorithms. The first approach involves the sequence of unitary operators alone, while the latter builds upon the former, utilizing an Oracle (black box) to replace key steps. Relying on an oracle occurs due to two potential reasons: firstly, the oracle might be trivial, and secondly, theory suggests that an efficient routine should exist to perform these steps, which are currently not a concern. In terms of the number of Oracle accesses, there is another way, quantum query complexity [33, 34], to analyze the complexity of a quantum algorithm.
Even though varied ways of describing quantum algorithms are commonplace in literature, this survey doesn’t strictly discriminate between them due to their inherent equivalency to one another. Still and all, we need to note that generating corresponding quantum circuits (gates) from unitary operators is not trivial. First, the popular superconducting quantum hardware has a sparse near-neighbor connectivity topology and needs to use swap gates to exchange non-adjacent qubits. How to layout the circuit, thereby reducing swap gates, is crucial to the performance. Second, the native gates supported by varied hardware are tremendously different; which gates are adopted to construct quantum circuits will also affect the performance. In fact, this issue, known as “quantum circuit compilation” is challenging and has been proven to be NP-hard [35].
Given that quantum circuits and quantum gates are fundamental for describing and running quantum algorithms, we will proceed to discuss the most frequently used elementary quantum gates in the subsequent section.
2) Quantum circuits and elementary quantum gates
Quantum algorithms can be viewed as unitary evolutions of a predetermined initial state within a quantum system. Keeping in mind that unitary operations are synonymous with quantum circuits, decomposing a quantum circuit into elementary gates essentially equates to segmenting complex unitary evolution into fundamental operations. A unified evaluation criterion, specifically a uniform set of elementary gates, streamlines the process of comparing and evaluating the efficiency and costs associated with various quantum algorithms. Conventionally, the elementary gates of quantum circuits are accounted for by single and double qubit gates. Importantly, the set of elementary gates is not static; a significant amount of variation can occur depending on the specific hardware support. Subsequently, we explore some commonly recognized elementary gates.
• single qubit gates
① Identity gate
$I=|0\rangle\langle 0|+|1\rangle\langle 1|$, i.e., the identity matrix. $I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
② Bit flip gate
$X=|0\rangle\langle 1|+|1\rangle\langle 0|$, $X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$.
When acting on $|0\rangle$, $|1\rangle$, and $|\phi\rangle=\alpha|0\rangle+\beta|1\rangle$:
$\begin{split}X&|0\rangle=|1\rangle\\ X&|1\rangle=|0\rangle\\ X&|\phi\rangle=X(\alpha|0\rangle+\beta|1\rangle)=\alpha|1\rangle+\beta|0\rangle\end{split}$ (21)
③ Phase flip gate
$Z=|0\rangle\langle 0|-|1\rangle\langle 1|$, $Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$.
When acting on $|0\rangle$, $|1\rangle$, and $|\phi\rangle=\alpha|0\rangle+\beta|1\rangle$:
$\begin{split}Z&|0\rangle=|0\rangle\\ Z&|1\rangle=-|1\rangle\\ Z&|\phi\rangle=Z(\alpha|0\rangle+\beta|1\rangle)=\alpha|0\rangle-\beta|1\rangle\end{split}$ (22)
④ Bit and phase flip gate
$Y=iXZ=i(|1\rangle\langle 0|-|0\rangle\langle 1|)$, $Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$
When acting on $|0\rangle$, $|1\rangle$, and $|\phi\rangle=\alpha|0\rangle+\beta|1\rangle$:
$\begin{split}Y&|0\rangle=i|1\rangle\\ Y&|1\rangle=-i|0\rangle\\ Y&|\phi\rangle=Y(\alpha|0\rangle+\beta|1\rangle)=i(\alpha|1\rangle-\beta|0\rangle)\end{split}$ (23)
X, Y, and Z are three important single qubit gates, also often represented as Pauli matrices: $\sigma^{x}$, $\sigma^{y}$, and $\sigma^{z}$. These three matrices are not only unitary but also Hermitian, and their eigenvectors form the coordinate axes on the Bloch sphere of a single qubit. They are often used to construct X-, Y-, and Z-axis rotation gates, which are the building blocks for parametric quantum circuits.
⑤ Rotation gates
$\begin{split}R_{x}(\theta)&=e^{-i\theta X/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}X\\ R_{y}(\theta)&=e^{-i\theta Y/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}Y\\ R_{z}(\theta)&=e^{-i\theta Z/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}Z\end{split}$ (24)
⑥ H gate
$H=\frac{1}{\sqrt{2}}\left((|0\rangle+|1\rangle)\langle 0|+(|0\rangle-|1\rangle)\langle 1|\right)$. Its matrix representation $H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$. When acting on $|0\rangle$, $|1\rangle$, and $|\phi\rangle=\alpha|0\rangle+\beta|1\rangle$:
$\begin{split}H|0\rangle&=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle)=|+\rangle\\ H|1\rangle&=\frac{1}{\sqrt{2}}(|0\rangle-|1\rangle)=|-\rangle\\ H|\phi\rangle&\!=\!H(\alpha|0\rangle\!+\!\beta|1\rangle)\!\!=\!\!\frac{1}{\sqrt{2}}(\alpha\!+\!\beta)|0\rangle\!+\!\frac{1}{\sqrt{2}}(\alpha\!-\!\beta)|1\rangle\end{split}$ (25)
$H$ gate has a broad application in generating quantum circuits. For example, acting on |$\ket{00\cdots 0}$ to generate the initial state for most quantum algorithms.
We need to note that the bit flip gate (X) is drawn differently in various literature (Fig. 2a and Fig. 2b), but they both represent the same thing.
Figure 2: The bit flip gate.
• Double qubit gates
Double qubit gates represent unitary operations encompassing two qubits. Prominent among them are the controlled unitary gates, where one qubit is the control end, and the other serves as the controlled qubit. The value of the control qubit, consequently, prescribes whether the unitary operation acts upon the controlled qubit.
In this scenario, we focus on the widely employed CNOT gate, which is alternatively known as the quantum Xor gate. This gate is associated with two inputs: |$\ket{x}$ and |$\ket{y}$, with |$\ket{x}$ in the control position and |$\ket{y}$ as the controlled entity. The semantics attached to the CNOT gate are as follows:
IF |$\ket{x}=\ket{0}$ THEN |$\ket{y}$ unchanged, perform I operation;
IF |$\ket{x}=\ket{1}$ THEN |$\ket{y}$ flipped, perform X operation;
Thereby, $CNOT:\ket{x,y}\rightarrow\ket{x,x\oplus y}$, the definition is:
$CNOT=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes X$ (26)
Acting effect:
$\begin{split}CNOT|00\rangle=|00\rangle\qquad CNOT|01\rangle=|01\rangle\\ CNOT|10\rangle=|11\rangle\qquad CNOT|11\rangle=|10\rangle\end{split}$ (27)
Matrix representation:
$\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$ (28)
Gate representation:
Figure 3: CNOT gate
• Treble qubit gate
It has been established that any unitary matrix within a $d$-dimensional Hilbert space can be decomposed into a product of two-level unitary matrices. Furthermore, it has been elucidated that a combination of single-qubit gates and a CNOT gate can facilitate any two-level unitary operation [36]. Consequently, without a doubt, any treble qubit gate can be simplified by employing a single-qubit gate and a dual-qubit gate. However, a unique treble qubit gate, known as the Toffoli gate, commonly features as a basic module within quantum circuits and usually doesn’t undergo further simplification:
The Toffoli gate comes with three inputs: |$\ket{x}$, |$\ket{y}$, and |$\ket{z}$, where the control qubits are |$\ket{x}$ and |$\ket{y}$, and |$\ket{z}$ operates as the controlled qubit. The semantics of the Toffoli gate are provided below:
IF |$\ket{x}=\ket{1}$ AND |$\ket{y}=\ket{1}$THEN
|$\ket{z}$ flipped, perform X operation;
ELSE
|$\ket{z}$ unchanged, perform I operation;
Therefore, $Toffli:|x,y,z\rangle\rightarrow|x,y,(x\land y)\oplus z\rangle$, the defination is:
$Toffli=|0\rangle\langle 0|\otimes I\otimes I+|1\rangle\langle 1|\otimes CNOT$ (29)
Acting effect:
$\begin{split}Toffli|000\rangle=|000\rangle\qquad Toffli|001\rangle=|001\rangle\\ Toffli|010\rangle=|010\rangle\qquad Toffli|011\rangle=|011\rangle\\ Toffli|100\rangle=|100\rangle\qquad Toffli|101\rangle=|101\rangle\\ Toffli|110\rangle=|111\rangle\qquad Toffli|111\rangle=|110\rangle\end{split}$ (30)
The matrix representation:
$\begin{pmatrix}1&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1\\ 0&0&0&0&0&0&1&0\end{pmatrix}$ (31)
Toffoli gate representation is shown in fig. 4a. In particular, when $|z\rangle=|0\rangle$, the Toffoli gate can be exploited as a quantum ‘AND’ gate (fig. 4b).
Figure 4: Toffli gate.
3) Measurement
Measurement (observation) is the only way to ascertain and extract the result of a quantum algorithm. From a physical implementation view, measurement embodies a degree of complexity. Observing or measuring implies an interaction between macroscopic or classical objects and a quantum system, with the process forming correlations and entanglements [37, 38]. Assume the state of the system under scrutiny to be: |$\ket{\phi}=\sum_{i}c_{i}\ket{i}$. The wave function of the entire system, together with the instrument, stands as following when the system is connected with the measuring apparatus:
|$\ket{\Psi}=\sum_{i}c_{i}\ket{i}\otimes e_{i}$ (32)
where {$e_{i}$} constitutes a set of the instrument’s wave functions. This correlation signifies a form of entanglement; once we detect that the instrument acquires the state |$\ket{e_{i}}$, the entire wave function collapses into the state |$\ket{i}\otimes\ket{e_{i}}$, and the system state |$\ket{i}$ can be inferred.
The physical realization of this process requires considerable experimental preparation, including tasks like calibrating the parameterized Hamiltonian, determining the ground state of the entire system, and so forth. Nonetheless, in this context, we focus solely on the algorithmic (circuit) level to utilize a simple example to illustrate the generation of entanglement, and the determining of the leftover segment when a partial state of an entangled system is observed.
Fig. 5 delineates the circuit employed to generate one of the Bell states - $\frac{1}{\sqrt{2}}\left(\ket{00}+\ket{11}\right)$. The first qubit |$\ket{0}$, transforms into a superposition state $\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right)$ following its passage through the $H$ gate. Subsequently, through the CNOT gate, the |$\ket{1}$ within the first qubit triggers a flip in the associated |$\ket{0}$ of the second qubit, thereby creating an entangled state (eq. 33). Post the generation of entanglement, these qubits are unable to be expressed as a $\otimes$ product of independent subsystems. When an observation of the second qubit is made and it is located in |$\ket{0}$, the first qubit collapses into |$\ket{0}$. Conversely, if it is in the state |$\ket{1}$, the first qubit also becomes |$\ket{1}$. In this operation, the first qubit can be conceived as the quantum system requiring observation, whereas the second qubit stands as the measuring instrument.
Figure 5: The circuit to generate Bell state
[…]
There are two possible outcomes of these observations, namely, $+1$ and $-1$, which correspond to the two eigenvalues of Pauli Z. The probability of obtaining $+1$ equals $|\alpha|^{2}$, whereas $|\beta|^{2}$ equates to the likelihood of garnering $-1$. The outcome of eq. 34 is the expected result of measurements. Specifically, if every observation yields a result of $+1$, it can be concluded that the qubit is in state |$\ket{0}$, and vice versa.
The aforementioned partial measurement on Bell state, whose observable can be written as $I\otimes Z$. Given a clear context, ’observed by $Z$’ can be simplified as measurements on the |$\ket{0}$ and |$\ket{1}$ basis (computational basis). It’s crucial to bear in mind the significance of the choice of the measurement basis. As an exemplification, for the state $\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right)$, using the |$\ket{0}$ and |$\ket{1}$ basis results in a $\frac{1}{2}$ probability for obtaining the post-measurement state |$\ket{0}$ and |$\ket{1}$, expressing the uncertainty of the result. However, if $\ket{+}$, $\ket{-}$ is employed (i.e., choosing observable Pauli $X$ for the measurement), the result $\ket{+}$ is guaranteed.
Being able to deterministically obtain the result of a quantum algorithm implies that the chosen measurement bases can reliably differentiate the final state. Although only orthogonal quantum states can assuredly be distinguished, the ultimate states of the quantum algorithm are occasionally non-orthogonal. In such instances, the objective is to ascertain the similarity rather than distinguishing the states reliably. This task can be proficiently executed by employing the subtle probability estimation of partial measurements. fig. 6 displays this technique, wherein the initial |$\ket{0}$ serves as an auxiliary qubit, transitioning into the superposition state $\frac{1}{\sqrt{2}}(\ket{0}+\ket{1})$ after passing through the left $H$ gate. Subsequently, under control of |$\ket{1}$ within this state, the circuit exchanges |$\ket{x}$ and |$\ket{y}$ to arrive at the final state eq. 35111The Swap gate can be implemented through the utilization of 3 CNOT gates.:
[…]
Figure 6: swap test to evaluate the similarity of two states
If measuring auxiliary qubit alone by |$\ket{0}$ and |$\ket{1}$ bases, the likelihood of acquiring the post-measurement state |$\ket{0_{anx}}$ is:
$P(\ket{0_{anx}})=\frac{1}{2}+\frac{1}{2}|\bra{x}\ket{y}|^{2}$ (36)
$|\bra{x}\ket{y}|$ | is referred to as fidelity in the field of quantum information, fundamentally akin to the classical cosine distance. Obviously, the probability equals 1 if |$\ket{x}$ and |$\ket{y}$ are entirely overlapped, while the probability is $\frac{1}{2}$ if |$\ket{x}$ and |$\ket{y}$ are orthogonal. It’s vital to emphasize that probability estimation doesn’t hinge on the dimension. The larger the dimension of the state space, the greater the efficiency of this quantum trick. Inferred from this, the general similarity measure, $Euclidean\ distance=\sqrt{2-2|\bra{x}\ket{y}|}$, can also be computed. In practice, this trick forms the foundational basis for myriad quantum machine learning algorithms, inclusive of supervised and unsupervised machine learning [39], kernel approaches in quantum machine learning [40, 41], Quantum Support Vector Machine (QSVM) [42], Quantum Nearest-Neighbor Algorithms [43], and so forth.
II-D The Description by Density Operators
The density operator is another more powerful mathematical language to describe quantum systems. The state vector can only represent pure states, while the density operator can represent pure and mixed states in a unified form. Besides, the density operator is a way more convenient tool to study quantum information, for example, the relationship between entanglement entropy and algorithm accuracy [44].
The density operator of a quantum system can be defined as follows. Suppose a quantum system is in pure state |$\ket{\psi_{i}}$ with probability $p_{i}$, where $i$ is the index of a set of state {|$\{\ket{\psi_{i}}\}$. Then, the quantum system can be described as an ensemble of pure states:
$\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$ | (37)
A density operator is also called a density matrix, and generally, these two terms can be substituted for each other without distinction. If the density matrix $\rho$ satisfies (1) $tr(\rho)=1$; (2) $\rho$ is a positive semidefinite matrix, then $\rho$ is a legal representation of an ensemble $\{p_{i},\ket{\psi_{i}}\}$.
With the help of the density operator, we can distinguish pure and mixed states. The pure state refers to the quantum system being in a deterministic state $\rho=\ket{\psi}\bra{\psi}$ |; that is, the quantum system is in |$\ket{\psi}$ with probability 1. In all other cases, the quantum system is in a mixed state. The criterion to judge whether a quantum system is in a pure state or a mixed state is as follows: if $tr(\rho^{2})=1$, $\rho$ is a pure state; If $tr(\rho^{2})<1$, then it is a mixed state. According to density operators, the quantum algorithmic key steps are described as: 1) unitary evolution
$\rho_{2}=U\rho_{1}U^{\dagger}$ (38)
where $\rho_{1}$ is the state in time $t_{1}$ and $\rho_{2}$ is the state in time $t_{2}$. 2) measurement
$\begin{split}p(m)=&tr(M_{m}M_{m}^{\dagger}\rho)\\ \rho_{m}=&\frac{M_{m}\rho M_{m}^{\dagger}}{tr(M_{m}M_{m}^{\dagger}\rho)}\end{split}$ (39)
where $\{M_{m}\}$ is a set of measurement operators, the index $m$ refers to the possible measurement outcomes. $p(m)$ is the the probability of obtaining the outcome $m$ and $\rho_{m}$ is the corresponding post-measurement state.
II-E QC in Noisy Intermediate-scale Quantum Era
The quantum computer was conceptually put forward in the early 1980s [24]. QC hardware has made remarkable progress such that it is possible to deploy quantum algorithms in real. Noisy Intermediate-Scale Quantum (NISQ) computing devices [45] demonstrates the quantum superiority over classical computing [46]. We are not going to cover some well-known algorithms, such as Shor’s algorithm [31] and Grover’s algorithm [47], since they have been over-rendered by numerous materials [14]. Another primary reason we don’t introduce these algorithms is that the corresponding circuit depths to implement these algorithms are far beyond what is achievable in the NISQ devices. In the NISQ era, the number of operations (quantum gates) to implement a quantum algorithm should remain moderate to alleviate the suffering from decoherence and quantum gate errors.
In this section, we focus on the promising apps that are suitable for NISQ devices, including Variational Quantum Eigensolver (VQE) [48], QAOA [49], and some quantum machine learning algorithms. These algorithms have a common characteristic: using hybrid classical and quantum iterations to calibrate parameterized ansatzes. We will describe how these algorithms work and also discuss the challenge in this category at the end.
1) VQE
VQE is mainly applied in computational chemistry. With the development of modern quantum chemistry, computational chemistry has become an important tool for drug synthesis, catalyst preparation, etc. However, chemical calculation involves a huge computation burden, whereas classical computers are stretched thin to provide adequate calculation accuracy and calculation scale. So, since quantum chemistry has to do with quantum systems, why not simulate computations on a quantum computer?
VQE simulates chemical computation on a NISQ quantum computer. It not only ensures the coherence of the quantum state but also achieves chemical accuracy for finding the ground state of a quantum chemical system. Its pipeline involves 3 major steps:
Step 1 – Pre-process
Firstly, a quantum chemical system (electron system of a molecular) is represented as a Hamiltonian by Born-Oppenheimer approximation:
$H=-\frac{h^{2}}{2m_{e}}\sum_{i}\nabla_{i}^{2}-\sum_{i}\frac{Ze^{2}}{r_{i}}+\sum_{i}\sum_{j}\frac{e^{2}}{r_{ij}}$ (40)
where $\frac{h^{2}}{2m_{e}}\sum_{i}\nabla_{i}^{2}$ is the kinetic energy term of the electron, $\sum_{i}\frac{Ze^{2}}{r_{i}}$ is the gravitational potential energy term between the nucleus and the electrons, and $\sum_{i}\sum_{j}\frac{e^{2}}{r_{ij}}$ is the repulsive energy term amongst electrons. This process is called the first quantization, and then $H$ in eq. 40 will be transformed to the structure represented by the fermion operators.
$H=\sum_{pq}h_{pq}a_{p}^{\dagger}a_{q}+\frac{1}{2}\sum_{pqrs}a_{p}^{\dagger}a_{q}^{\dagger}a_{r}a_{s}$ (41)
where $a_{j}^{\dagger}$ and $a_{j}$ are known as creation and annihilation operators, respectively. This process is called second quantization. After this process, the number of qubits to characterize $H$ will drop from scale $\mathcal{O}(mlog(n))$ to $\mathcal{O}(n)$ [50, 51], where $m$ and $n$ are the number of electrons and basis functions, respectively.
Finally, it needs to translate the fermionic Hamiltonian into Pauli operators. Fermionic Hamiltonian obeys the antisymmetry relationship, thereby the linear combination of translated Pauli strings should maintain this property. Exploiting Jordan-Wigner transformation [52], Parity transformation [53], Bravyi–Kitaev [53, 54, 55], or other improved methods [56], the $H$ composed by Pauli operators satisfying this criterion can eventually be obtained.
$H=\sum_{j}w_{j}H_{j}$ (42)
where $w_{j}$ are the weights of the Pauli strings in Hamiltonian.
Step 2 – Core-process
The core step is sketched in Fig. 7. First, the trail state is initialized as the Hartree-Fock wavefunction. Simply speaking, for the molecular system with electronic spin-orbits, the Hartree-Fock state can be represented as |$\ket{0\dots 01\dots 1}$, where |$\ket{0}$ denotes empty orbit and |$\ket{1}$ denotes occupied orbit. Then, this state goes into the parameterized ansatzes (circuits) for evolving. All the parameters $\vec{\theta}$ (in Fig. 7) of the ansatzes could be initialized randomly or by some specific methods to alleviate the suffering from barren plateaus. When the trial wave function |$\ket{\varphi({\vec{\theta}})}$ is obtained, it must be rotated into the Z basis (computational basis) for measurements. In terms of the expectation value on each component (eq. 43), we can get the total expectation $\langle{H}\rangle$ by weighted summation eq. 42. Based on the $\langle{H}\rangle$ computed, we update the ansatz parameters using a certain optimization strategy and begin a new iteration until obtaining $\min\limits_{{{\vec{\theta}}}}\langle{H}\rangle$.
$\langle{H_{j}}\rangle=\bra{\varphi({\vec{\theta}})}U_{j}^{\dagger}(Z^{\otimes n})U_{j}\ket{\varphi({\vec{\theta}})}$ (43)
Figure 7: The schematic diagram of VQE core-process.
Step 3 – Post-process
The post-process, i.e., error mitigation, is an additional layer of processing on the measurement results aimed at reducing the impact of quantum noise for more accurate results. Since this thing is involved in all practical QC programs, we present a unified discussion after introducing other NISQ algorithms.
2) QAOA
The QAOA [49] is mainly used to solve combinatorial optimization problems. Due to the ubiquitousness of optimization in various scientific and technological domains, this algorithm has aroused broad research interests. The QAOA is an approximate version of the Quantum Adiabatic Algorithm (QAA) [57]. By applying the Suzuki-Trotter theorem, the time-dependent Hamiltonians of QAA are broken into a sequence of $2p$ time-independent Hamiltonians and induce two parameterized unitary operators $\hat{U}_{B}(\beta)=e^{-i\beta\hat{B}}$ and $\hat{U}_{C}(\gamma)=e^{-i\gamma\hat{C}}$, where $\hat{B}$ and $\hat{C}$ are the initial Hamiltonian and final Hamiltonian of QAA corresponding to “mixing Hamiltonian” and “problem Hamiltonian” of QAOA, respectively. Acting these two unitary operators with tuneable parameters $\beta$ and $\gamma$ on the initial state iteratively, the final state |$\ket{\beta,\gamma}$ will approximate the optimal solutions to the optimization problem.
|$\ket{\beta,\gamma}=\hat{U}_{B}(\beta_{p})\hat{U}_{C}(\gamma_{p})\dots\hat{U}_{B}(\beta_{1})\hat{U}_{C}(\gamma_{1})\ket{s},$ (44)
The schematic diagram of QAOA shown in fig. 8, it involves 3 major steps:
Figure 8: The schematic diagram of QAOA.
Step 1 – Prepare initial state
In the vanilla QAOA for solving an unconstrained problem – MAXCUT, the initial state is prepared as $\ket{++\dots+}$, i.e., the uniform superposition over all possible states. When QAOA is applied to solve constrained problems, the straightforward way is to add penalty terms into the problem Hamiltonian to ensure infeasible solutions that violate the constraints are inferior to optimal feasible solutions when calculating the objective function. This scheme, say the Quadratic Unconstrained Binary Optimization (QUBO), does not modify the original operator framework, therefore can still use $\ket{++\dots+}$ as its initial state.
Despite this, numerous variant QAOA-based algorithms have shown that the choice of initial states can be more than $\ket{++\dots+}$. The naive approach is if a pre-processing method can obtain a biased initial solution (state) close to the optimal solution, then the convergence toward the final state should reach faster. This so-called “warm-starting” approach takes a wide variety of techniques, including variable relaxation, positive semi-definite programming, graph neural network, matrix product state, etc. to this end.
The other initial state preparation is for the alternative approach for handling constraints. That approach aims at restraining the evolution in the feasible subspace, thus $\ket{++\dots+}$ including infeasible states beyond the range of the feasible subspace is not applicable for the commencement. In this case, the initial state should be prepared as a superposition state composed of all or part of feasible solutions satisfying the constraints and excluding infeasible solutions. For a general constrained optimization problem, such as the Maximum Independent Set, preparing a uniform superposition state composed of all feasible solutions is equivalent to preparing an arbitrary uniform superposition state, which costs a lot. (exponential cost for deterministic manner [58] or post-measurement selection in a probability manner [59]). But this approach works particularly well for a certain class of NP-hard problems – Graph partition [60], maximum k-vertex cover [61], max-k colorable subgraph [62] etc., as their feasible subspace is spanned by all permutations of bit-strings with $Hamming\ weights=k$ (The number of $1s$ in a bit string is equal to $k$). That means their optima-searching space can be represented by a specific form of Dicke states – |$\ket{D^{n}_{k}}=\frac{1}{\sqrt{\tbinom{n}{k}}}\sum_{x\in\{0,1\}^{n},wt(x)=k}\ket{x}$, which can be easily prepared in deterministic manner [63].
Step 2 – Construct mixing Hamiltonian and problem Hamiltonian
As the name suggests, the mixing Hamiltonian $\hat{B}$ offers “mixing” operations. If regarding $\beta$ as a time-dependent parameter, $e^{-i\beta\hat{B}}$ actually describes a continuous quantum walk on the graph $\hat{B}$, on which the walker can reach any connected nodes from the initial state (node). In other words, the mixing operation is a specific unitary operation evolving the initial state to a superposition state mixed with other feasible states. In the scheme for unconstrained problems, or the QUBO-based scheme for constrained problems, the mixing Hamiltonian is always defined as $\sum_{i}\sigma^{x}_{i}$. In physics, this operator is called the transverse field operator, and its corresponding graph is a hypergraph, which connects pairwise nodes by $Hamming\ distance=1$.
In the alternative scheme for tackling constrained problems, mixing Hamiltonian designing is more subtly, since eliminating penalty terms and restricting the dynamics to the feasible space can get potential benefits of faster convergence with lower cost. Hadfield et al. proposed the seminal approach – Quantum Alternating Operator Ansatz, in which they constructed XY-mixer, Partial mixer, etc., for varied constraint types [64]. In recent two years, there are emerging many variants mixing Hamiltonian [65], which aims at evolving the initial state to the target state through smaller $p$ iterations via the mixing ability. A smaller $p$ means a shallower quantum circuit, thereby more efficiency and less suffering from decoherence.
Problem Hamiltonian $\hat{C}$ is used to encode the objective function, i.e. $\bra{x}\hat{C}\ket{x}=C(x)$, where $x$ is a bit string whose components “0/1” indicates the corresponding combinatorial variable is included in the solution or not; $C(x)\in\mathbb{R}$ is the objective function under the variable $x$. Depending on the optimization problem at hand, the highest/lowest energy state |$\ket{x^{*}}$ of Ising Hamiltonian $\hat{C}$ encodes the optimal solution. For Ansatz-based schemes, one needs to formulate each component of $x$ with $\frac{I-\sigma_{i}^{z}}{2}$ (the so-called spin to binary relation) and then use it to generate matrix $\hat{C}$; for QUBO-based schemes, still needs to construct extra penalty terms in $\hat{C}$, the tricks may refer to [66].
Step 3 – Generate the corresponding circuit on specific hardware and run
In terms of $\hat{U}_{B}$ and $\hat{U}_{C}$, one translates these operators to circuits for execution. The initial state goes through the circuit, gets the final state |$\ket{x}$, does measurement $\bra{x}\hat{C}\ket{x}$, and then calibrates the values of $\vec{\beta}$ and $\vec{\gamma}$ for the next cycle until |$\ket{x}$ converges to |$\ket{x^{*}}$. This procedure, like VQE, also involves the choice of classical strategies and error migration methods, which we will discuss in the subsequent sections.
Besides, note that implementing a QAOA-based algorithm in a real QC device is more challenging since the operator form of its variant versions is more complex. Moreover, current QC devices differ widely in qubits-connecting topology, native gates set, fidelity, etc., and compiling software offered by vendors is far from being able to abstract these underlying physical characteristics. Thus, a lot of additional hard work needs to be done to achieve good performance [35].
3) Quantum Machine Learning
Quantum machine learning, with which physicists have been fascinated for more than two decades [67], holds a solid foundation since the matrix manipulating techniques used in AI are well-studied in quantum theory. For this reason, and with the AI wave, people hope to further fuel the surge in AI research through the rich theory and technique of quantum technology. And this is indeed the case, many remarkable achievements have emerged in recent years [68]. For example, QSVM has a logarithmic runtime in both training and classification stages, achieving an exponential acceleration compared with the classical counterpart [42]. Nevertheless, the current capabilities of QC devices are lagging behind since its core subroutine – quantum principal component analysis (QPCA) [69] employs quantum Fourier transform, in which the controlled unitary gate is decomposed into a deep quantum circuit far beyond the capacity of QC devices in NISQ era. Given there is a long way toward large-scale fault-tolerant QC, quantum machine learning and other quantum algorithms require finding ways to adapt to the limited scale of imperfect QC resources [70]. That means the quantum algorithms running on the shallow-depth quantum circuit could still achieve some practical benefit. To this end, the hybrid workflow – uses parameterized quantum circuits (PQC) to run on a quantum computer and then optimizes parameters on a classical computer, becomes a typical QNN architecture. To some extent, PQC is the synonym of QNNs, and VQE and QAOA could be regarded as two special cases of PQC [71]. The quantum machine learning literature is so rich that our discussion is not going to cover all research branches. Instead, we focus on some hot topics: classification, generative AI, and RL.
• Classification
Classification, broadly also including clustering, is the most studied topic in machine learning. In PQC-based frameworks, quantum subroutines replace one or several steps in a classical neural network aims at improving the neural network’s abilities of expressibility, trainability, generalization, etc. [71], which Llyod calls quantum embedding [72]. Fig. 9 shows a typical architecture of such frameworks. The leftmost block denotes a preprocessing step on the classical computer to extract feature data. This step can be simple as, for example, in terms of the gray value of an image pixel, embedding it as the rotation angle of the rotation gate in eq. (45) for the subsequent quantum processing [73], where $a_{i}\in[0,255]$ is the gray value of a pixel..
$R_{y}(\theta_{i}),\quad\theta_{i}=\frac{a_{i}}{255}*\pi/2.$ (45)
Still, it can become more complex. [74, 75] employ quantum autoencoders to extract distinguishing characteristics from quantum states in a classical latent space, so as to use reduced quantum data for clustering and classification. Quantum embedded circuits for encoding is usually neither optimized nor trained, which would affect the computational power of QNNs [76].
Figure 9: The schematic diagram of typical hybrid quantum convolutional neural network architectures.
The abstract notation $U(\phi)$ and $U(\theta)$ in the middle quantum processing block represent two unitary operations for mapping classical data to the quantum domain and generating entanglement, respectively. $U(\phi)$ and $U(\theta)$ could be varied a lot in terms of different scenarios. In quantum convolutional neural networks (QCNNs), $U(\phi)$ and $U(\theta)$ are combined together as a quantum convolution layer implemented by a cascade of several layers of $R_{y}(\theta)$ rotation gate and CNOT gates [73]. Another version of QCNNs combines the multi-scale entanglement renormalization ansatz (MERA) to implement a pooling layer besides the quantum convolutional layer [77]. One point we would like to emphasize is that entanglement is essential to achieve a quantum advantage using hybrid neural networks. In a quantum transformer architecture, the entanglement is generated by RBS gates, which can be decomposed further by four $Hadamard$ gates, two $R_{y}$ rotation gates, and two two-qubit $CZ$ gates [78, 73].
After quantum data are remeasured, they become classical data again to continue the remaining operations in classical neural networks. Depending on the problem at hand, varied objective/loss functions are computed, and the result is used to calibrate the variational values $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ until convergence. This process involves training iteratively and dealing with noise and gate errors that all PQC-based algorithms confront. Therefore, we will discuss this in a subsequent independent section.
• Generative AI
In recent years, there is no doubt that the hottest AI technology is generative AI, which generates new samples by taking the underlying probability distribution of a dataset such as images, videos, etc. From the Boltzmann machine [79], to the latest Generative Adversarial Networks (GAN) [80] and Denoising Diffusion Probabilistic Model (DDPM) [81], all have corresponding quantum versions.
The quantum generative learning models (QGLMs) seem to be more versatile than their classical counterparts. Besides image generation [82], there are works discuss quantum state estimation [83], continuous distributions generation [84], and distribution learning [79] etc., show their ability for accomplishing what the classical tools are incapable of.
QGLMs adopt the same workflow as their classical counterparts. For example, the Quantum QGLMs (QuDDPM) [81], shown in Fig. 10, is also composed of forward noisy and backward denoising process. The forward noisy diffusion process exploits a quantum scrambling circuit that evolves the sample data towards a random ensemble of pure states by applying a series of random unitary gates independently. On the other hand, the backward denoising process reduces noise gradually towards the final generated data by performing projective measurements after applying unitaries on ancilla qubits. This process ensures that the generated data maintains purity and mimics the sample data. The integration of QC techniques and diffusion models offers benefits that are not only suitable for efficient training and learning of quantum data but also improve performance in classical tasks. As an example, [85] show two quantum diffusion models – Q-Dense and QU-Net outperform classical counterparts in terms of performance metrics like FID, SSIM, and PSNR.
Figure 10: The schematic diagram of QuDDPM [81].
The key challenge in QGLMs is exploring their learnability in terms of generalization, expressivity, and trainability. The generalization and expressivity of QGLMs are less studied compared to quantum discriminative learning models. The definition of generalization in quantum generative learning is vague under different model settings, making it difficult to study the generalization of QGLMs. The trainability of QGLMs is also challenging due to the non-convex loss landscape and the presence of barren plateaus. The performance of QGLMs can be significantly affected by quantum hardware imperfections such as gate noise and measurement error. When the level of noise exceeds a threshold, the power of near-term quantum computers may be inferior to classical computers. Quantum system noise can also lead to diverged optimization in QGLMs.
• RL
Given that QC can enhance computational power, thus naturally, combining QC and RL was conceived and has been studied for more than a decade [86]. QRL can be roughly divided into three categories according to the proportion of ‘quantum’: the first is so-called inspired RL, which basically belongs to the category of classical heuristics. The second is PQC-based RL, which can run on NISQ devices and will be the focus of our subsequent discussion. The last one is full-quantum RL. As the name suggests, algorithms in this category only work well on fault-tolerated quantum computers.
PQC-based RL can be further classified in terms of the role of PQC in the pipeline, one is the value-function approximation approaches; one is policy approximation approaches, and the combined approximation for value and policy, i.e., actor-critic (AC) approaches. A typical PQC-based QRL pipeline is summarized in Fig. 11. The hybrid agent observes the initial environment and applies some pre-processing to encode the observation into the feature map $U(\phi)$. Through a $\theta$-parameterized unitary operation $U(\theta)$, the current state |$\ket{s_{t}}=U(\theta)U(\phi)\ket{0}^{\otimes n}$ can be gotten. Then depending on the measurements of the action-selection observable $\bra{s_{t}}O_{a}\ket{s_{t}}$, the state-action value function $Q_{\theta}(s_{t},a_{t})$ or the policy $\pi_{\theta}(a_{t}|s_{t})$ will navigate the next favorable action $a_{t+1}$. Finally, the cumulative reward $r_{t}$ computed by interacting with the environment is iterated into a classical optimizer for consecutive parameter updates and optimization until convergence. The overall iterative pipeline is similar to the classical DRL, except that state change occurs in the quantum domain, where the variational unitaries could be regarded as a kind of QNNs for approximation.
Figure 11: The schematic diagram of PQC-based RL.
Applying quantum approaches to RL presents several challenges. One key challenge is the limited exploration of the true potential of near-term quantum approaches in RL, as existing works have struggled to solve classical benchmarking tasks using PQCs. Another challenge is achieving sufficient performance in RL environments when using PQCs as value-function or policy approximators, as demonstrated in recent works. Additionally, the complexity and resource-intensive nature of quantum computations pose challenges in terms of scalability and practical implementation in RL tasks. Furthermore, the interpretability of quantum learning models is a concern, as the quantum nature of computations make these models difficult to understand compared to classical ones. Lastly, the accessibility of quantum computers for training QRL models potentially exacerbates existing inequalities in AI development and usage.
4) Challenges and Current Solutions
PQC-based algorithms, or rather hybrid Variational Quantum Algorithms (VQA), all have a subroutine that employs classical optimizers to enable the training of parametrized quantum circuits to find the optimality of the objective function. However, barren plateaus become a core issue that confines the training efficiency. In the following section, we will summarize the related work regarding the circumstances where barren plateaus may arise and the solutions to avoid/alleviate them. On the other hand, due to the underlying discrepancy of QC infrastructure and various types of noise, running quantum algorithms on a genuine NISQ device is far more complex than running them on a simulator, which is another challenge that will be discussed afterward.
• Barren plateaus: Barren plateaus (BP) are characterized by an exponential vanishing of gradients across the optimization landscape. These plateaus not only impede gradient-based optimizers but also pose challenges for gradient-free algorithms like Nelder-Mead and COBYLA, as the flat landscape tends to exponentially dampen the cost function [87]. The emergence of BP limits the potential of PQCs to achieve quantum advantages. This is because the exponential decline of gradients necessitates exponential resources for optimization, whereas quantum algorithms typically aim for polynomial scaling of the system. Consequently, as problem sizes increase, PQCs become increasingly difficult to train.
BP occurs in varied cases: random initialization in parametric circuit [88], objective function defined in terms of global observables [89], entanglement-induced [90], noise-induced [91], and PQC structure-induced [92], etc. For various scenarios, people have come up with methods to alleviate or avoid the occurrence of BP. Layerwise learning, transfer learning, and Bayesian learning have been used to find a favorable initial set of parameters for mitigating BPs [93]. [94] proposed an algorithm called generalized quantum assisted simulator, which does not require a classical-quantum feedback loop, to bypass the BP problem. [95] established a fundamental connection between expressibility and trainability by extending the concept of barren plateaus to arbitrary ansatz, which provides insights into designing strategies for problem-inspired ansatz to avoid BPs. [96] show that over-parametrization in QNNs can mitigate BPs.
Note that although there has been a wealth of research on the BP problem, little is known about the optimization landscape of a generalized PQC-based algorithm. To the best of our knowledge, the only one that can be sure to avoid BP is the work [97]. The paper rigorously analyzes the gradient scaling for parameters in QCNNs architecture. It finds that the variance of the gradient vanishes no faster than polynomials, indicating that QCNNs do not exhibit barren plateaus. This result is significant as it provides an analytical guarantee for the trainability of randomly initialized QCNNs.
• Run quantum algorithms in real quantum computers: As the first step, the programmer must decompose unitary operators (matrix) in the quantum algorithm into small ones – single-qubit and double-qubit gates. For a generalized unitary matrix $e^{-iH}$, till now, we only know that Hamiltonian $H$ must be sparse, or low-rank [69], then it can have a polynomial decomposition. For decades, numerous related works have struggled to find methods for more efficient decomposition. Next, the programmer translates logical gates into native gates that the specific hardware supports. Since the underlying supports for current QC vary widely in technology (superconducting, ion trap, optical photon, etc.), and even with the same technology, they still differ in low level, a lot of extra work needs to be done for the optimization [98]. For example, IBM and Rigetti ’s machines are both grounded in superconducting technology. The former has the native CNOT gate, while the latter needs to translate $CNOT(A,B)$ gate into a sequence of rotations and CZ operations: $R_{z}(\pi/2)B;R_{x}(\pi/2)B;R_{z}(\pi/2)B;CZ(A,B);R_{z}(\pi/2)B;\\ R_{x}(\pi/2)B;R_{z}(\pi/2)B$ [99]. Compilers provided by vendors, such as IBM’s Qiskit, can translate multi-control gates into native gates customized for their architecture.
Finally, programmers need to find ways to mitigate various types of errors. In the NISQ era, hardware resources cannot support the significant consumption of applying quantum error correction code. Quantum error mitigation, which suppresses errors via classical postprocessing techniques, is the only practical approach to handle errors at this stage[70]. It requires the noise model in advance and then estimates parameters for the expected value of the noise-free model by analyzing how the noise model behaves under different noise levels. The other is probabilistic error cancellation based on gate set tomography [100], which does not know the precise physical noise model in advance. It combines linearly independent basis set operations and gate set tomography to fully remove the impact of localized Markovian errors by systematically measuring the effect of errors. This technique has been successfully implemented in trapped-ion systems and superconducting systems.
II-F Quantum-inspired Algorithms
The origin of quantum-inspired algorithms can be attributed to the integration of QC concepts into conventional algorithms. Typically, algorithms in this class are described as running on classical computers, with some steps imitating properties unique to the quantum realm.
A subclass of these algorithms mentions the implementation on a quantum computer. Consider a QRL algorithm, which denotes the state transition via the measurements of an observable [86]. The previous state will collapse to one of the observable eigenstates (next state) associated with an action chosen probability by the agent. In theory, this seems plausible. Nevertheless, for a complex application scenario, physically representing and implementing an observable at each step could be an issue. Moreover, the no-cloning principle makes it unclear how to obtain sufficient independent quantum state copies for the statics of measurement results. Therefore, the algorithm is classified as quantum-inspired RL [101].
The second subclass of these algorithms explicitly states that they are running classically. A well-known example is Tang’s work [102], in which a binary tree-like data structure is proposed to satisfy quantum state preparation assumptions. A classical algorithm exploiting this structure can output an $\mathcal{l}^{2}$-norm sample from a rank-k approximation of a matrix in nearly polynomial time. The third category is named quantum-inspired metaheuristic algorithm. As the name suggests, these algorithms modify steps of classical metaheuristic algorithms inspired by quantum principles so as to improve performance [103]. There are many related research works, including quantum ant colony algorithm [104], quantum evolutionary algorithm [105], QGA [106], QPSO [107], to solve well-known fundamental NP-hard problems such as knapsack, traveling salesman to practical issues involving clustering, project scheduling, etc.
The modifications to quantum-inspired algorithms based on QC principles follow a similar pattern: utilizing qubits, or n-dimensional normalized vectors, to map and represent the original data, such as chromosomes in genetic algorithms; employing rotation gates or matrices to evolve qubits for exploration; using a random number generator to generate new feasible solutions, simulating quantum measurements collapsing to ground states; integrating entangled states in qubit representation, so that when creating a new population by measuring the qubit string (solution), the entangled qubit pairs in the solution diversify the data and help in avoiding stagnation of local optima.
Quantum-inspired metaheuristic algorithms enhance optimization performance by leveraging QC concepts to accelerate convergence, enhance exploration and exploitation, and find the best solutions more effectively. However, most published work still requires in-depth complexity analysis to justify performance benefits. On the other hand, how to appropriately embed QC into optimization workflow, be compatible with the metaheuristic algorithms, and not reduce the algorithm’s speed is still a concern.
II-G Common QC Platforms and Libraries
1) Hardware
• Superconducting: Superconducting is the most popular technical route to produce QC hardware. This technique uses superconducting circuits to generate and manipulate qubits and employs microwave pulses to control quantum gates. IBM Quantum and Google Sycamore quantum computers are both grounded in this technique route.
-Pros: fast response, scalable, and seamlessly integrated with existing microelectronics technology. Moreover, these circuits can be mass-produced and integrated at the chip level.
-Cons: Superconducting qubits are susceptible to environmental perturbations, including thermal noise, electromagnetic noise, etc., resulting in shorter coherent time.
-Challenges: maintain high fidelity while scaling.
• Ion Trap: Ion trap QC is a relatively mature technique that traps ions with light and magnetic fields and manipulates them with lasers. Typical products include IonQ Aria and Honeywell System Model H.
-Pros: High stability, long coherent time, and high quantum gate fidelity, making them well-suited for QC. Moreover, the ion trap technique can practically manipulate all qubits, unlike the superconducting technique, which requires a long chain of operations on non-adjacent qubits.
-Cons: Cumbersome physical manipulation for ion traps, high optical stability requirements, and complex laser operating systems.
-Challenges: complex to scale.
• Optical QC: Using photons as information carriers. Xanadu Borealis is one representative machine.
-Pros: less affected by environmental noise and have relatively mature techniques for generating, manipulating, and detecting single photons.
-Cons: require precise optical operations, reliable single-photon sources, and efficient single-photon detectors.
Challenges: implement quantum logic gates with high efficiency and high fidelity, as well as large-scale optical QC.
• Topological QC: Topological QC relies on emergent quasiparticles, called Majorana zero-modes, that are robust against perturbations. Microsoft has been engaged in this technique route.
-Pros: more stable qubits, resistant to many types of error and noise.
-Cons: manufacturing and controlling Majorana zero-modes is complex and technically challenging.
Challenges: have not reliably fabricated and detected Majorana zero-modes in experiments.
2) Software
• Qiskit [108]: Qiskit is an open-source software development kit (SDK) developed by IBM. It provides tools for programming on IBM quantum computers or simulators on a local computer. Qiskit uses the Python programming language. Its official documentation provides a wealth of tutorials and case studies for QC practitioners to learn, practice, and conduct research. It is worth mentioning that Qiskit, combined with IBM’s quantum computers, provides an excellent QC ecosystem for researchers worldwide. As a software project, Qiskit has developed rapidly and constantly integrated new achievements. For example, Qiskit Machine Learning has been incorporated into the Qiskit platform, providing various tools and algorithms for building and training quantum machine learning models, which can handle numerous tasks such as classification, regression, and RL.
• Cirq & TensorFlow Quantum [109]: Cirq is developed by the Google AI Quantum Team. It provides a powerful simulator and also supports programming for specific quantum processors. It has two distinctive components: OpenFermion, mainly applied to quantum chemistry, and TensorFlow Quantum (TFQ), an open-source machine learning library based on TensorFlow. It introduces two primitive data types: Quantum circuit batch, which represents the quantum circuit with varied sizes, and Pauli sum batch, which represents a linear combination of tensor products of Pauli operators. Exploiting these two data types, TFQ can efficiently sample the circuit batch’s output distribution, compute the expected values of Pauli sum and circuit batches, calculate the gradient, and simulate quantum circuits and quantum state batches. TFQ practitioners can not only leverage TensorFlow’s functionality, such as automatic differentiation and machine learning tools, but can also utilize quantum circuit simulators and connect quantum hardware. Yet, since TFQ is closely related to TensorFlow and Cirq, it faces a compatible challenge with other QC libraries or frameworks.
• PennyLane [110]: PennyLane is a Python-based quantum machine learning library that seamlessly integrates classical machine learning algorithms with QC using automatic differentiation. It operates independently of any specific underlying framework and smoothly interfaces with both classical machine learning frameworks (like PyTorch and TensorFlow) and QC frameworks (such as Qiskit and Cirq). Utilizing a plugin system, PennyLane enables users to connect to various quantum hardware and simulators without the need for custom code for each platform. Its libraries and modules encompass Quantum-specific operations for evaluating and optimizing quantum circuits, Quantum nodes for encapsulating quantum circuits usable in machine learning, plugins facilitating the execution of quantum circuits on real hardware or simulators, and Quantum Machine Learning components for integrating quantum circuits with classical machine learning models, including neural networks, to develop hybrid quantum-classical machine learning models. PennyLane is highly cross-platform compatible, has automatic differentiation quantum nodes for training machine learning models, and provides user-friendly interfaces for quantum algorithms and applications that are easy to understand and implement. On the other hand, its performance and functionality strongly depend on the stability and quality of the plugin system. Besides, the PennyLane community is relatively small and needs more supporting resources than other tools grounded in traditional machine learning frameworks.
• Torch Quantum (QuantumNAS) [111]: Torch Quantum, developed by a consortium including MIT, Yale, and the University of Texas at Austin, integrates seamlessly with PyTorch, adding several key components: a Quantum State Vector Simulator optimized for GPU acceleration, Quantum Circuit Modules enabling the definition and manipulation of PQCs akin to classical neural networks, Quantum Optimization facilitating parameter optimization using classical optimization algorithms like gradient descent, and an Interface to real Quantum devices. This extension of PyTorch allows developers to construct quantum programs using familiar tools and interfaces while benefiting from automatic differentiation for training QNNs. Additionally, the High-Performance Computing State Vector Simulator supports large-scale simulations on GPUs. However, it may not match the performance of specialized QC platforms like IBM Quantum and Rigetti due to its reliance on PyTorch optimization.
III QC in Communications
III-A Overview
This section explores the applications of QC in various aspects of wireless communication, including power allocation, channel assignment, and user association, as illustrated in Fig. 12. By leveraging the principles of quantum mechanics, researchers aim to address the challenges faced by traditional communication systems and unlock new opportunities for optimizing network performance and enhancing user experience.
Figure 12: Different network scenarios in wireless communications.
III-B Power Allocation
III-B1 Overview
Power allocation within wireless communication [112] involves the strategic distribution of available power resources among the diverse components of a communication system, such as base stations and user devices. The primary objective is to optimize power utilization to achieve communication that is both reliable and efficient, taking into account factors such as signal quality, network coverage, interference management, and energy consumption [113], which is also an important issue in 5G/6G cellular networks [114]. In more straightforward terms, power allocation revolves around determining the appropriate amount of power assigned to each transmitter or device in a wireless network to ensure effective communication.
Power allocation is usually implemented at the physical layer, involving only nodes within one same communication cell, optimizing the transmission power of signals to maximize transmission efficiency or minimize bit error rate. For example, power optimization allocation schemes with relay selection have been studied [115]. By selecting appropriate relay nodes and allocating corresponding power, the energy efficiency is maximized. Power allocation in wireless communication systems is primarily aimed at mitigating issues arising from the considerable variation in received signal strength due to varying distances between transmitters and receivers. This variance can lead to degradation in system performance, necessitating power control to balance the transmission power.
TABLE III: Summary of QC Works on Power Allocations in Wireless Networks
QC category Network structure Work Network details Highlights Optimization objectives QC alg. Control manner
Conventional QC Single cell communication [116] K transmitter-receiver pairs with Gaussian Interference Utilize a QC approach to solving power allocation Maximize weighted-sum rate QPSO Centralized
[117] Primary users and secondary users competing for spectral resources Employ Bernstein approximation to address probability constraints and co-evolutionary theory to handle large-scale problems. Maximize network throughput
[115] One-hop cooperative relay between BS and multiple users in a NOMA system Apply QPSO to maximize QoS from the perspective of MIMO Maximize computation efficiency and energy efficiency
Overlay networks [118] Primary users and one primary BS; Cognitive users and one cognitive BS Formulate resource scheduling in a two-layer wireless network, and propose a model to describe outage probability Minimize system outage probability
[119] K users and N sub-carries in OFDMA system: two heterogeneous services, minimum-rate guarantee and best effort services Divide the problem: channel assignment given power allocation, power allocation given channel assignment; transform it into a concave optimization and solved by Dinkelbach method Maximize energy efficiency
AI-driven QC Single cell communication [120] Multiple users and a BS in a NOMA system Propose federated teleportation based QNN and implement it in a case study Maximize the weighted-sum rate QNN Distributed
[121] MIMO in a NOMA system Present a QNN with parallel training Maximize the weighted-sum rate
[122] Users and a BS in a NOMA system Present quantum circuits of layerwise embedding with implementation in NISQ computers Maximize cumulative energy efficiency over time QDRL Centralized (only one UAV)
TABLE IV: Details of QC Algorithms for Power Allocation
Work Algorithm Orth. QC Implemented in quantum computers Platform No. qubits
[116] QPSO Computer simulation 0
[117] QPSO Computer simulation 0
[115] QPSO Computer simulation 0
[118] QPSO Computer simulation 0
[119] QPSO Computer simulation 0
[120] QNN IBM Qiskit 16
[121] QNN IBM Qiskit 9
[122] QDRL IBM Qiskit 12
In this section, we undertake an investigation of existing QC-driven approaches along two fundamental dimensions: the complexity of network scenarios and the categories of QC algorithms. The former encompasses networks ranging from one-hop NOMA/OFDMA systems to multiple-cell and overlay systems. Meanwhile, we classify the latter into conventional QC algorithms, such as classical QAOA and quantum-inspired QPSO, and AI-driven QC algorithms like QNN, with an emphasis on comprehending their suitability and effectiveness across various network landscapes.
III-B2 Applications of Conventional QC for Power Allocation
QC in a single-cell system:
A single-cell communication system in wireless networks typically consists of a BS or access point and multiple user devices, responsible for handling wireless communications within a local area. User devices within this area communicate with the BS. A typical single-cell system is shown in Fig. 12. [116] and [117] considered power allocation within a straightforward context, focusing on a single cell of a multiple access (MA) system while applying conventional QC algorithms. $K$ transmitter-receiver communication pairs with Gaussian interference were involved in [116]. The problem was conceptualized as a non-convex and NP-hard programming formulation aimed at maximizing the weighted-sum rate. To address this problem, a QPSO based approach was employed, leveraging the characteristics of QC and biological behaviors. However, the network structure was simple and only Shannon capacity was considered without addressing interference. Energy limitation issues was addressed in [117] in cognitive radio sensor networks by constructing a mathematical model for power allocation. Bernstein approximation technique was used to control the approximation error of function value to tackle probability constraints. In addition, it adopted the concept of co-evolution to deal with large scale problems. The proposed optimization problem was complex and NP-hard, which was corresponding to the co-evolutionary theory. A stochastic optimization algorithm based on polynomial approximation and QPSO was proposed to solve the problem.
Though one-hop cooperative relay between the BS and multiple users, such as UE 1 communicated with BS via relay-UE 2 in Fig. 12, was considered [115], it was still within the area of a single-cell system. The relay link was introduced between the source node and destination node to reduce channel interference. Thus, there were two different links: the BS to relay nodes, and relay nodes to users. The strategies of power allocation for multiple relays and users significantly influenced the interference among nodes in the local communication area. The situation became worse when a full-duplex operation of channels from the perspective of MIMO was considered. Besides, another QoS metric, secrecy rates of end-users, was considered to represent the overhearing from eavesdroppers near end-users. The problem formulation was proved to be non-convex and solved by QPSO.
QC in an overlay system:
In contrast, more complex scenarios, spanning from multiple-cell networks to overlay networks, were investigated in [118], and [119] while still utilizing conventional QC algorithms. An overlay system refers to an architecture in which additional network components or services are deployed over a single-cell system, like mesh networks, or virtualized network functions, and among others.
The problem of scheduling multiple resources including power allocation and channel assignment in a overlay OFDMA network was studied in [118]. There was an additional constraint on the channel assignment for cognitive users, which complicated the problem comparing with resource scheduling in traditional wireless cognitive networks without primary networks. The work formulated the problem of resource scheduling in a two-layer wireless network, and proposed a model to describe system outage probability according to strategies of power allocation and channel assignment. [119] discussed the resource scheduling, including power allocation and channel assignment, in a single-cell OFDMA system with $K$ users and $N$ sub-carries, in which there were two heterogeneous service requirements from users: minimum-rate guarantee services and best effort services. It can be viewed as a virtual overlay network according to two heterogeneous service requirements. The work divided the problem formulation into two optimization parts: channel assignment for a given power allocation, which was solved stochastically by QPSO based solution; power allocation for a given channel assignment, which was transformed into a concave optimization problem and solved optimally by Dinkelbach method. Finally, an acceptable solution is obtained by iterating two optimization parts. Noting that there is no analysis to obtain the optimal solution by iteration of two part even if the optimal solution is obtained.
III-B3 Applications of AI-driven QC for Power Allocation
The previous studies on power allocation rely on conventional QC algorithms like QPSO. However, it is worth noting that these conventional algorithms still demand significant computational resources and lead to high overhead, such as computation delays. Recent research has turned towards AI-driven QC solutions for power allocation. These approaches, such as QNN and QDRL, leverage advanced machine learning techniques to mitigate the computational complexities inherent in traditional QC algorithms. By harnessing the AI power, these methods offer promising avenues for more efficient and effective power allocation in wireless networks. For example, [120] proposed a distributed QNN framework for the purpose of optimizing wireless communication and gave a case study in non-orthogonal multiple access system for the framework implementation. The objective was also to maximize the weighted-sum rate under consideration of channel interference, and the proposed QNN algorithm was implemented in quantum circuits, which accelerates running time overhead.
Other works implemented in quantum computers include [121, 122] in the field of power allocation. [121] presented a QNN with parallel training for optimizing wireless resource allocation. The proposal reduced the dimension of training data by sending only the statistical parameters of the dataset instead of the whole dataset. The proposed method was applied to optimize transmit precoding and power allocation in MIMO-NOMA. The analysis showed that the parallel training approach achieved a comparable sum rate with lower complexity compared to conventional methods. It highlighted the potential benefits of QNNs in processing large dimensional solution spaces and addressed the challenges of distributed datasets and training complexity. [122] proposed a layerwise QDRL method for optimizing continuous large space and time series problems using deep layer training. It leveraged the advantages of QC to maximize reward and reduce training loss. The study focused on jointly optimizing UAV trajectory planning, user grouping, and power allocation for higher energy efficiency. The main contributions of the study were the introduction of layerwise quantum embedding, and the presentation of quantum circuits for practical implementation. However, it was assumed that users can perceive perfect CSI in the UAV downlink communication channel, which limits its application.
In conclusion, the integration of QC techniques in power allocation holds promise for optimizing various aspects of wireless networks. Table III provides an overview of the main findings, optimization objectives, and QC algorithms utilized across different wireless networks. Noting that, there are relatively fewer studies employing QC algorithms for power allocation, largely due to the localized nature of power allocation in wireless networks, typically confined to a single-cell system. This observation is reflected in both the choice of QC algorithms and their control manners as depicted in Table IV. The majority of studies utilize QC simulated algorithms that are implemented on traditional computers.
III-C Channel Assignment
III-C1 Overview
Channel assignment is a fundamental aspect of wireless communication network management, where the task involves allocating communication channels to different users or devices within the network and well addressed in 5G networks [114], such as slicing solutions. In wireless networks, communication channels represent the available frequency bands or time slots that facilitate data transmission between transmitters and receivers. Efficient channel assignment is crucial for optimizing network performance, minimizing interference, and maximizing spectral efficiency. The problem has been addressed by various algorithms and strategies, such as RL approaches [123]. These methods aim to allocate channels in a way that minimizes interference between users while maximizing the utilization of available spectrum resources.
With the growing complexity and scale of wireless networks, traditional channel assignment approaches may face limitations in addressing the dynamic and heterogeneous communication environments. QC presents a promising avenue for tackling these challenges by offering novel algorithms and techniques that can potentially improve the spectral efficiency, reliability, and scalability of channel assignment in wireless communication networks.
TABLE V: Summary of QC Works on Channel Assignment in Wireless Networks
QC category Network structure Work Network details Highlights Optimization objectives QC alg. Control manner
Conventional QC Single-cell systems [124] K users and a receiver over MIMO antennas Encode the maximum likelihood detection problem into a Hamiltonian operator, and transform the problem into a level-p QAOA circuit Maximize signal detection probability QAOA Centralized
[125] N transmitter-receiver pairs in a cell Transform sub-channel assignment into a capacitated Max k-Cut formulation, which is induced as a QUBO problem, and QA is applied to solve the problem Maximize the sum rate of D2D transmitter-receiver pairs QA Centralized
[126] N users competing for orthogonal channels in cognitive networks Map channel assignment matrix to the chromosome of the algorithms to decrease search space Maximize sum rates of users, the minimal rate of users, proportional fair rates of users QGA Centralized
[127] Mobile cellular networks Present a case study on the assignment problem of root sequence index in LTE/NR physical random access channel configuration Maximize orthogonality metric, the set cardinality of root sequence index, among mobile users QA Centralized
Multi-hop networks [128] No two links within k hops is simultaneously activated. Schedule air-links for maximum throughput while avoiding k-hop interference Maximize network throughput QA Centralized
[129] Network graph: vertex denotes T/R nodes and edges denote links. Schedule air-links for minimizing k-hop interference Minimize k-hop interference QA Centralized
AI-driven QC Single-cell systems [130] NOMA networks Propose a QNN framework for resource allocation and a solution for non-orthogonal NOMA as a case study Maximize the sum communication rate of users. QNNs Centralized
[131] Multiple UAVs a central server Study the impact of channel characteristics, such as communication scattering. Formulate UAV movement into MDP, and develop a quantum multi-agent AC algorithm Maximize sum communication rates of ground users Quantum multi-agent AC Distributed
Overlay networks [132] Vehicle-to-everything: eNodeBs allocate channels to RSUs, RSUs allocate them to vehicles. Formulate channel assignment as a discrete MDP and adopt QRL to find the optimal policy. Minimize the transmission delay and energy consumption QRL based on Grover Search Algorithm Distributed
Conventional QC for IRS systems Single-cell systems [133] One BS and multiple users in a cell under assistance of one IRS Propose a scheduling method where the IRS is assigned to only a single user at a given time slot. Maximize network throughput QA Centralized
[134] One BS and users in a cell under assistance of one IRS Propose a scheduling method where IRS is assigned to only one user at a given time slot. Maximize secrecy energy efficiency Quantum bald eagle search algorithm Centralized
[135] One BS and users in a cell with assistance of multiple IRSs Propose a scheduling method where one IRS is assigned to only one user at a given time slot Maximize network throughput QA Centralized
Multi-cell systems [136] Multiple BSs and users with assistance of multiple IRSs Propose a scheduling method where each IRS is assigned to only one user at a given time slot Maximize network throughput QA Centralized
[137] Multiple BSs and multiple users under assistance of one IRS Share the IRS among BSs by periodically switching phase shifts to reduce hardware cost for multiple IRSs Minimize the number of required time slots to complete the communication QA Centralized
TABLE VI: Details of QC Algorithms for Channel Assignment
Work Algorithm Orth. QC Implemented in quantum computers Platform No. qubits
[124] QAOA IBM Qiskit 1, 2, 3
[125] QA D-Wave 2000+
[128] QA D-Wave 164, 405
[129] QA D-Wave 2X 164, 405, 759, 783
[127] QA D-Wave No. optimization variables
[126] QGA Computer simulation 0
[132] QRL Computer simulation 0
[131] QDRL Computer simulation 0
[130] QNNs, RL-QNN IBM Qiskit $12(N_{\text{layer}}-1)N_{\text{neuron}}+$ 1, $13+3(N_{\text{layer}}-1)N_{\text{neuron}}+$ 1
[134] QGA Computer simulation 0
[135] QA D-Wave $N_{\text{user}}N_{\text{IRS}}N_{\text{time slots}}$
[133] QA unknown unknown unknown unknown
[136] QA D-Wave unknown
III-C2 Applications of Conventional QC for Channel Assignment
QC in a single-cell system:
The application of QAOA was discussed to solve the problem of maximum likelihood detection of binary symbols transmitted over MIMO channel [124]. The maximum likelihood detection problem is NP-hard and difficult to solve using classical computers. The study utilized QAOA to encode maximum likelihood detection into a level-p QAOA circuit, which was solved using classical optimizers. The work provided an overview of QAOA and their applications in solving various problems.
D2D communications with $N$ transmitter-receiver pairs in a single-cell system was discussed in [125]. The authors highlighted the challenge of mutual interference between D2D and nearby cellular communication and the need to reduce the interference through resource allocation including subchannel assignment and power allocation. They formulated the problem as a capacitated Max k-Cut and applied QA. Specifically, sub-channels were assigned to disjoint transmitter groupings, and the issue was formulated as a QUBO problem. The algorithm was implemented in a quantum computer. Though the paper considered a simple scenario of subchannel allocation, it provided guidelines to apply optimization techniques of QC to solve similar problems. Applying another QC algorithm QGA for cognitive radio spectrum allocation, illustrated in Fig. 12, in a single-cell network was discussed in [126]. Specifically, the work proposed a mapping process between the channel assignment matrix and the chromosome of the algorithms to decrease the search space. Thought it highlighted the importance of dynamic spectrum access in order to improve spectrum utilization in the validation results, there is no consideration in the problem formulation and analysis.
The preceding works explored QC algorithms for channel assignment in local wireless networks. Their application in mobile cellular networks was also conducted in [127]. It presented a case study on the root sequence index assignment problem in LTE/NR physical random access channel configuration. They formulated the problem as a QUBO problem and solved it using a cloud-based QC platform. The results showed that QA solver can successfully assign conflict-free root sequence index, but comparison with classical heuristics revealed that some classic algorithms were more effective in terms of solution quality and computation time. The reason lies in the semi-quantum algorithm and the type of quantum computer, which provides reference for the correct use of QC techniques.
QC in relaying networks:
The problem of channel assignment in relaying networks is typically formulated as conflict graphs that were resolved by QA in [128] and [128]. A real-world application of QA, specifically the scheduling of air-links for maximum throughput while avoiding interference near network nodes was studied in [128]. It introduced a novel gap expansion technique that benefits QA more than simulated annealing, leading to improved speedup and network queue occupancy. The proposed QA was implemented in quantum computers. However, it is assumed that overall network information was given in the scheduling algorithm. [129] explored network performance in the same scenario with [128]. Weighted maximum independent set was utilized to model the K-hop interference. It made a comparison between simulated annealing and QA in solving the scheduling problem. The results showed that the QA solution outperformed simulated annealing when using the gap expansion technique.
III-C3 Applications of AI-driven QC for Channel Assignment
QC in a single-cell system:
A framework of QNNs for resource allocation [130] was designed in wireless communications, such as 5G/6G, to reduce time complexity while maintaining performance, and a reinforcement-learning-inspired QNN was proposed to further improve performance. Specifically, they presented the quantum circuit design of the QNN to ensure practical implementation in NISQ computers. The QNN was applied to group users in NOMA as a case study. The performance on channel assignment outperformed other solutions, but at the expense of a large number of qubits. RL based QC algorithms, quantum multi-agent AC networks, were performed in [131] to construct a robust mobile access system with multiple UAVs. The algorithm utilized MARL to enable UAVs to learn collectively and optimize their actions within a shared environment. The study addressed scalability challenges by implementing a quantum centralized critic. One of its advantages lied in the design of a realistic environment with noise injection, the evaluation of QMARL and noise reflection on training policies.
One of 6G scenarios, Metaverse, was discussed in [132]. The problem lied in insufficient sampling in model training for connected and autonomous vehicles (CAVs). To efficiently achieve Metaverse, channel resources among RSUs and vehicles were significant for network performance, in which eNodeBs allocated channel resources to RSUs and RSUs allocated them to vehicles. It formulated the channel resource assignment as a discrete Markov Decision Process (MDP) and adopted QRL to obtain the optimal policy. However, the study assumed that vehicle information was given without considering their dynamic, such as moving out communication ranges of RSUs.
III-C4 Applications of QC Algorithms in IRS
Intelligent Reflective Surfaces (IRSs) [138] represent a revolutionary technology in 6G communications that leverages passive elements to manipulate electromagnetic waves. These surfaces are composed of programmable materials capable of adjusting the amplitude, phase, and polarization of incident electromagnetic signals as illustrated in Fig. 13a. By intelligently controlling these properties, IRSs can effectively shape and redirect wireless signals to enhance communication performance.
(a) One BS and one IRS
(b) One BS and multiple IRSs
(c) Multiple BSs and multiple IRSs
(d) Multiple BSs and one IRS
Figure 13: Different IRS systems
[135] proposed an approach to controlling the propagation path characteristics from the transmitter to the receiver by appropriately according the reflection phase of IRS elements in a communication cell consisting of one BS and multiple users under assistance of multiple IRSs as illustrated in Fig. 13b. The core problem was to schedule the assignment of IRSs to a given users for a unit time. The authors first formulated the scheduling problem as a QUBO problem and then developed an algorithm based on QA to solve the issue. The solution was validated in the quantum computers with the platform D-Wave. There are other works [133] and [134] on the same scenario as [135]. Taking [133] as an example, it compared its tensor contraction and regression based QA with PSO and factorization machine based QA. However, there was no explanation for its advantage over others.
[135] and [133] considered the scheduling problem with only one BS in a communication cell. To improve system scalability, [136] tackled the IRS sub-channel assignment with multiple BSs as illustrated in Fig. 13c. The network structure was different from the scenario in which there was only one BS, when multiple BSs were introduced into the network. In this work, IRSs and users were coordinated by a BS and a server (such as a cellular station), but were not necessarily in the same coverage cell of the BS. When there were multiple BSs in different cells under assistance of multiple IRSs, the channel scheduling problem was formulated as a combinatorial optimization, which was also in the area of QUBO. And a QA based solution was proposed to solve the problem. However, it was a centralized scheduling method for multiple IRSs. In addition, there was no detailed discussion on QA implementation though the simulations validated its effectiveness.
[135] and [133] considered network performance in a communication cell with multiple users, an IRS and one BS. The network was easily extended to a communication cell with multiple users, multiple IRSs, and multiple BSs [136] in different cells. To reduce the network overhead such as hardware costs, it obtained the similar network performance with [136] by sharing an ISR of switching phase shift as illustrated in Fig. 13d, which was studied in [137]. Specifically, when a particular phase shift was set at a given time, the corresponding time resources and BSs were assigned to UEs existing in the reflection direction of the radio waves transmitted from each BS, which enabled efficient communication. In the case of multiple operators, each operator performed such radio resource-allocation control in parallel, enabling communications with the IRS without coordination required between the operators. The process was formulated as QUBO and solved by classical QC algorithm. However, it can not achieve the optimal performance since they worked independently. And thus, it is still viewed as a centralized approach. Other works related to IRS scheduling in the area of channel assignment includes [134] that considered both power allocation and communication security. Its scenario was similar with [135] and [133] in a communication cell. However, the solution for channel assignment of scheduling phase shift was independent with other optimization variables. The quantum-inspired bald eagle search algorithm was developed to solve the problem, which was not orthodox QC.
In summary, QC-enabled algorithms offer a promising solution to enhance performance in channel assignment, as the overview of these studies presented in Table V and QC details in Table VI.
III-D User Association
III-D1 Overview
User association in wireless networks refers to the process of determining which access point, BS, or sensor sink that a user device connects to for communication. This decision is crucial for optimizing network performance, ensuring efficient resource utilization and service quality [139]. The primary goal of user association is to provide seamless connectivity while maximizing network capacity and throughput. Various factors influence user association decision, including signal strength, available bandwidth, network load, user mobility, and QoS requirements. Note that network dynamics, particularly, sink (such as UAVs) mobility, are one of key features in user association, which was explored in numerous studies [140], such as in mobile sensor networks and UAV networks. Solutions encounter computational constraints like with a substantial number of devices and dynamic environment. The challenge can be addressed through the utilization of QC techniques. Thus, in this section, we primarily categorize our discussion according to two key criteria: QC algorithm category and mobility of BSs or sinks.
III-D2 Applications of Conventional QC for User Association
Without consideration of mobility:
TABLE VII: Summary of QC Works on User Association in Wireless Networks
QC category Network dynamic Work Network details Highlights Optimization objectives QC alg. Control manner
Conventional QC Without mobility [141] Multiple sensors communicate with a BS under assistance of cluster head Propose a quantum-inspired evolutionary algorithm to select clusters from multiple sensors Maximize network lifespan Quantum-inspired evolutionary algorithm Centralized
[142] Multiple sensors communicate with a BS under assistance of cluster head Propose a quantum-inspired gravitational search algorithm to select clusters from multiple sensors Maximize network lifespan Quantum-inspired gravitational search algorithm Centralized
[143] Multiple sensors communicate with a BS under assistance of cluster head Propose a quantum-inspired gravitational search algorithm to assign operational states for multiple sensors Maximize network lifespan Quantum-inspired gravitational search algorithm Centralized
[144] Multiple sensors communicate with a BS under assistance of cluster head Propose a quantum elite gray wolf optimization to assign operational states for multiple sensors Maximize network multiple QoS Quantum elite gray wolf optimization Centralized
[145] Multiple sensors communicate with a BS under assistance of cluster head Propose a Quantum Clone Whale Optimization Algorithm to assign clusters from multiple sensors Maximize network multiple QoS Quantum clone whale optimization algorithm Centralized
[146] Multiple sensors communicate with a BS under assistance of cluster head Propose a Quantum Artificial Bee Colony to assign clusters from multiple sensors Maximize network lifespan Quantum artificial bee colony Centralized
[147] Multiple sensors communicate with a BS under assistance of cluster head Propose a QPSO to assign clusters from multiple sensors Maximize network lifespan QPSO Centralized
[148] Multiple sensors communicate with a BS under assistance of cluster head Propose a QGA to assign clusters from multiple sensors under assistance of cluster head Maximize network lifespan QGA Centralized
[149] Multiple sensors communicate with a BS under assistance of cluster head Propose a QPSO to assign clusters from multiple sensors Maximize network lifespan QPSO Centralized
[150] Online social networks Predict online social connectivity to recommend user connection Predict link connectivity Quantum ant colony optimization Centralized
Mobility [151] Wireless sensor networks with a mobile sink Analyze path planning of a mobile sink and transform the problem to a travelling salesman problem Find the shortest path of a mobile sink QA Centralized
[152] Multiple-UAV networks Organize multiple-UAV into clusters while considering three metrics namely average distance, distance to UAVs and the number of UAV neighbors Optimize the fitness function with average distance, distance to UAVs and the number of UAV neighbors QGA Centralized
[153] Multiple-UAV networks to provide computation resources dynamically to ground users Propose a distributed leader-follower coalition formation algorithm in a multiple-UAV networks to accomplish service provision dynamically in an unknown environment Multiple objectives of minimizing coalition cost, maximizing reliability and reputation QGA Distributed
AI-driven QC Mobility [154] Multiple UAVs communicate with multiple BSs, and there is only communication between a UAV and a BS Propose a distributed leader-follower coalition formation algorithm in a multiple-UAV networks to accomplish service provision dynamically in an unknown environment Minimize the movement penalty and connection outage duration QDRL Distributed
[155] Multiple ground users communicate with a UAV Optimize UAV trajectory planning, NOMA user grouping, and power allocation, solved by proposed QDRL that can tackle issues of a large and continuous action and state spaces Maximize communications throughput per unit of energy consumed by the UAV QDRL Centralized
The study examineda basic wireless sensor network [141], consisting of sensors and a BS. It introduced a clustering method based on QEA to extend the network lifespan. However, it does not account for network dynamics, such as traffic patterns and device mobility. In [142], a comparable scenario to [141] was explored, with the introduction of the quantum-inspired gravitational search to enhance energy efficiency and prolong network lifespan. The study considered communication constraints and application-specific criteria, encompassing parameters such as mean relative deviation, spatial density error, and sensor-out-of-range error. Nonetheless, the paper lacked an in-depth analysis of the optimality of the proposed solution. Meanwhile, [143] presented an alternative quantum-inspired approach to the issue addressed in [141] and [142]. Named the quantum-inspired gravitational search algorithm, this solution aimed to determine the operational states of sensors while preserving network connectivity. It introduced a weighted fitness function, integrating connectivity, energy, and application-specific parameters. However, accurately determining the appropriate weights for these factors posed a challenge.
Several additional quantum-inspired approaches have been proposed in the literature, including works such as [144, 145, 146, 147, 148, 149]. These contributions introduced techniques such as quantum elite gray wolf optimization, quantum clone whale optimization, quantum artificial bee colony, QPSO, and QGA, all aimed at addressing the same problem outlined in [141]. Additionally, other studies explored link connectivity within QC algorithms. For example, [150] investigated online social networks like Facebook, Weibo, and Twitter. The goal was to predict link connectivity, facilitating recommendations for registered users to connect with others.
With consideration of mobility:
The preceding examination of user association in the absence of accounting for mobility and network dynamics has been extensively explored in existing literature. Subsequent studies consider more complex scenarios involving UAV networks and wireless sensor networks featuring mobile sinks. For example, [151] focused on optimizing the path planning of a mobile sink to enhance energy efficiency. In this scenario, the mobile sink started its path from a designated location, traversed through the communication range of each sensor along the predetermined path, conducted a round of data collection, and subsequently returned to its initial starting point. They employed mathematical transformations to align the problem characteristics with those of a traveling salesman (TSP) problem. An optimal TSP path was obtained by using quantum tunneling effect and quantum circuit to achieve parallelism.
Regarding UAV networks, [152] considered a scenario with multiple UAVs but did not address data relay among UAVs. The research designed an objective function incorporating parameters like average distance, distance to neighboring UAVs, and the number of UAV neighbors to select cluster heads. While the proposal accounted for UAV mobility, it lacked mobility analysis within the model. Additionally, the paper explored UAV clustering application in image classification, but the model lacked a direct correlation between UAV clustering and image classification. In contrast, [153] tackled the challenge of UAV coalition formation for providing computation resources in a large-scale, heterogeneous UAV network. The study operated in an unknown and dynamic environment, where UAVs must dynamically acquire environmental knowledge. Leader UAVs assessed follower UAVs’ contributions in resource availability after completing missions, sharing this information during coalition formation. QGA guided the inclusion of idle UAVs into coalitions, considering UAV mobility characterized by randomness and lacking correlation with ground points of interest (PoI). However, the study lacked an analysis of PoI impact on system performance, differing significantly from edge computing research, particularly in task offloading investigations [156].
III-D3 Applications of AI-driven QC for User Association
Several other studies [154][155] integrated AI techniques with QC algorithms within their solutions while considering network dynamics. For example, [154] discussed the path planning for cellular-connected UAVs to minimize the weighted sum of time cost and expected outage duration. Traditional offline optimization techniques face challenges due to practical considerations like local building distribution and directional antenna radiation patterns. To overcome this, the authors proposed a DRL solution that mapped the problem into a MDP, enabling the UAV to find the optimal flying direction within each time slot. However, the analysis did not establish direct correlations among multiple UAVs, simplifying the model into individual path planning for each UAV.
Due to robust capacity of QDRL for addressing challenges including those characterized by a large state and action spaces, it can also be effectively applied in the optimization of multiple objectives. [155] addressed multiple interconnected challenges, encompassing UAV trajectory planning, NOMA user grouping, and power allocation. Within this framework, users within a group maintained communication with the UAV within their communication ranges along the UAV trajectory. The decision regarding energy consumption by the UAV directly influenced the objective. Given the extensive and continuous nature of both action and state spaces, the paper introduced a QDRL approach to effectively tackle the complexity of the problem.
In conclusion, QC algorithms present a compelling approach for user association in wireless networks. Table VII offered a consolidated overview of these studies. However, there are two open issues in these QC algorithms addressing user association. First, these works typically relied on computer simulations for QC algorithm implementation. Second, few studies accounted for the impact of mobility on performance, even in UAV networks.
IV QC in Networking
IV-A Overview
Figure 14: Different network scenarios in wireless networking.
This section explores how QC principles can revolutionize various aspects of networking, including routing, task offloading, and content caching as illustrated in Fig. 14. We aim to provide valuable insights into the current state of research, challenges, and future directions in leveraging QC to enhance networking capabilities.
IV-B Routing
IV-B1 Overview
Routing presents fundamental challenges in network systems, encompassing the task of determining optimal paths for data packets to travel from a source to a destination. In wireless networks, where devices may be mobile and network conditions dynamically change, routing becomes even more complex. The primary goal of routing is to ensure efficient and reliable data transmission while minimizing latency, packet loss, and resource consumption. Existing approaches to routing encompass a variety of algorithms and protocols designed to address different aspects of routing optimization. Traditional methods like Optimized Link State Routing Protocol (OLSR) and Ad hoc On-Demand Distance Vector (AODV) attempt to dynamically establish and update routes based on network topology changes. However, they suffer from overhead due to frequent updates and incur latency in route discovery. Additionally, traditional routing algorithms may not adequately address challenges such as mobility, variable channel conditions, and energy constraints in wireless networks [157, 158].
QC-driven approaches offer promising solutions to overcome the limitations of traditional routing methods in wireless networks. By leveraging the principles of quantum mechanics, they can explore vast solution spaces more efficiently and potentially find optimal or near-optimal routes in significantly less time. In addition, QC enables the development of AI-driven routing algorithms that can adapt and learn from network dynamics, optimizing routing decisions in real-time. These quantum-enhanced approaches have the potential to improve routing efficiency, minimize latency, and enhance network reliability in wireless communication systems.
TABLE VIII: Summary of QC Works on Routing in Wireless Networks
QC category Network feature Work Network structure Highlights Optimization objectives QC alg. Control manner
Conventional QC Without consideration of mobility [159] Wireless mesh networks with multi-hop relays in a 6G scenario The paper gives a detailed example of applying QAOA in solving a single-objective formulation. Maximize the network utility including battery cost, sum throughput for all nodes, and the number of hops QAOA Centralized
[160] A fleet of vehicles with a central depot is routed to visit a set of customers Propose a multi-phase method using Ising machines to minimize travel distance while balancing depot load and vehicle load. Find the shortest travel route while balancing loads. QA Centralized
[161] A fully connected multi-hop network with a single source node, a single destination node and a cloud of relay nodes Develop a multi-objective dynamic programming framework for generating Pareto-optimal routes relying on the correlations of the specific links Identify the entire set of weakly Pareto-optimal routes with multiple objectives Boyer-Brassard-Hoyer-tapp Quantum Search Algorithm Centralized
[162] Self-organizing network in which data are relayed to destinations by intermediate nodes Propose an optimal multi-objective quantum algorithm to evaluate paths using Pareto optimality Maximize network utility: overall bit error rates, linear-domain path losses and delays Boyer-Brassard-Hoyer-tapp Quantum Search Algorithm Centralized
[163] Wireless sensor networks with multi-hop relays Propose a QC solution focusing on quantum representation, measurement, and rotation angle Minimize the total energy cost over all relaying paths Quantum ant colony optimization Centralized
[164] Wireless mesh networks Propose a quantum ant colony algorithm to predict pheromone values to measure paths Predict pheromone values Quantum ant colony optimization Centralized
[165] Mesh networks in IoT Improve the robustness of network connectivity for IoT topology Sum of the number of nodes in maximum connected sub-graph QGA Centralized
[166] Wireless sensor networks relaying data by multiple hops to sinks Propose a QGA for multi-sink deployment while balancing energy consumption among sensors Maximize network utility including the number of hops, energy consumption, and sink dispersion QGA Centralized
[167] A weighted digraph wireless mesh network: source and destination nodes as multicast trees QoS multicast routing is converted into ILP with QoS constraints, solved by QPSO combined with loop deletion operation Minimize network delay QPSO Centralized
[168] A directed network graph in industrial IoT Study the impact of p-degrees of network connectivity on network performance Minimize energy consumption and network delay QPSO Centralized
[169] Cooperative data transmission from a cluster head to BS Propose a QPSO algorithm to assign relay-cluster heads Maximize network lifespan QPSO Centralized
Consideration of mobility [170] Mobile ad-hoc networks Embed Q-learning of RL in quantum genetic strategy Minimize local energy consumption QGA Centralized
[171] Mobile ad-hoc networks Embed augmented Q-Learning algorithm and combine the OLSR to optimize the selection of multi-point delay sets Maximize the network utility: overhead of topology control, delivery packet rates and time delay of packet transmission QGA Centralized
[172] UAV path planning Combine the differential evolution algorithm with QPSO Sum of the required time to finish the tasks on each target QPSO Centralized
[173] Mobile multi-hop networks with communications from UAVs to BSs under relay assistance Design a method for neuro-fuzzy cluster based topology construction and find routes for inter-cluster UAV communication Minimize delay and power consumption Quantum ant lion optimization Centralized
AI-driven QC Without consideration of mobility [174] Wireless sensor networks with multi-hop relays from the source nodes and destination nodes A QRL algorithm is developed based on Grover iteration for efficient decision-making Maximize the network utility of energy, channel and data buffer QRL Centralized
[175] Multi-robot network: autonomous mobile robots to identify locations and detect surrounding static and dynamic objects Propose a centralized training and decentralized execution QMARL for multi-robot control and coordination Maximize network utility of transportation delay, load balancing, and the ratio of positive delivering goods QMARL Distributed
Consideration of mobility [176] Multiple UAVs as mobile BSs to gather data from other devices Propose a QMARL for massive agents Reward maximization QMARL Distributed
TABLE IX: Details of QC Algorithms in Routing
Work Alg. Orth. QC Quantum computer Platform No. qubits Mobility
[159] QAOA IBM QC 65
[160] QA Fixstars Amplify AE, GUROBI, SA, Christofides, and Feld Unknown
[161] Boyer-Brassard-Hoyer-tapp Quantum Search Algorithm Unknown Unknown
[162] Boyer-Brassard-Hoyer-tapp Quantum Search Algorithm Unknown Unknown
[163] Quantum ant colony optimization Computer simulation 0
[164] Quantum ant colony optimization Computer simulation 0
[165] QGA Unknown Unknown
[166] QGA Computer simulation 0
[167] QPSO Computer simulation 0
[168] QPSO Computer simulation 0
[169] QPSO Computer simulation 0
[170] QGA Computer simulation 0
[171] QGA Computer simulation 0
[172] QPSO Computer simulation 0
[173] Evolutionary Quantum Pareto Optimization Computer simulation 0
[174] QRL Quantum virtual machine 4
[175] QMARL Computer simulation 0
[176] QMARL VQC Number of UAVs
IV-B2 Applications of Conventional QC for Routing
Without consideration of mobility
The paper [160] introduced a multi-phase method to address the vehicle routing with backhauls, mapping key elements of each phase onto QUBO models suitable for Ising machines. The two-phase method comprised a clustering phase followed by a routing phase, while the three-phase method included a depot clustering phase, a vehicle clustering phase, and a routing phase. The clustering-related phases were expansions of the knapsack problem, while the routing phases were tackled as TSP. Its process of modeling the problem as a QUBO formation can be regarded as a tutorial for applying QC in the field of networking.
The paper [161] presented a multi-objective dynamic programming framework aimed at generating potentially Pareto-optimal routes by leveraging correlations among specific links within these routes, thereby significantly reducing the total number of routes. This framework extended the widely single-objective Bellman-Ford algorithm to handle multiple objectives. Additionally, the paper introduced a quantum-assisted algorithm that capitalized on both the dynamic programming framework and the synergies between Quantum Programming and Hybrid Programming to effectively solve the multi-objective routing problem. Similar works from the authors with different objectives or different QC algorithms were presented in [177] and [178]. The study [162], similar to [161], utilized the Boyer-Brassard-Hoyer-Tapp quantum search algorithm, but in a different scenario within self-organizing networks. They introduced an optimal multi-objective quantum-assisted algorithm aimed at assessing legitimate routes while incorporating the concept of Pareto optimality to reduce complexity. However, the study overlooked the influence of node mobility, a critical characteristic in self-organizing networks like Ad-hoc networks.
The energy optimization model introduced in [163] targets sensor-enabled IoT environments, deriving energy consumption as the cost of energy-oriented paths. A QC-driven solution was proposed for optimizing the problem, emphasizing quantum representation, measurement, and rotation angle. However, a notable drawback of this approach was that minimizing the total energy cost across all relaying paths did not necessarily equate to prolonging the network lifetime. There were other works [164] based on quantum ant colony algorithms to solve routing problems. It utilized the metric, pheromone value inspired by ant colony algorithm, to measure a routing path. Due to dynamic network changes, it required a large computation time to track pheromone traits, especially in a large wireless sensor network.
Additional popular quantum evolution based proposals for routing issues included QGA solutions [165, 166] and QPSO solutions [167, 168, 169]. [165] improved the network robustness of connectivity for IoT topology duo to node failures caused by malicious attacks. It proposed a QGA algorithm to maximize the cumulative sum of the number of nodes in the maximum connected sub-graph of the network. Additionally, [166] was designed to be resilient, which adapted to any changes in the planned topology of deployed sinks in anticipation of sensor failures. Also it achieved energy consumption balance among sensors by equalizing the path length and brings down the cost of wireless sensor networks by minimizing the number of sinks. Multicast routing in wireless mesh networks were considered with the objective of minimizing network delay [167]. The routing problem was converted into an integer programming problem with QoS constraints and was solved by QPSO combined with loop deletion operation. Connectivity in routing was explored in [168, 169]. [168] considered heterogeneous scenarios of network topology for optimal path configuration by exploring and exploiting the hunts. It applied multiple inputs from heterogeneous industrial IoT into QPSO techniques. The differential evolution operator and crossover operations were used for information interchange among the nodes to avoid trapping into local minima. Specifically, it studied the impact of p-degrees of network connectivity on network performance. The same scenario of IoT was also studied in [169] to optimize selection of cluster heads for multi-hop inter-cluster communication, which was different from user association in a single-cell communication in Subsection III-D. It proposed a cooperative multiple-input-multiple-output connectivity scheme between sensors and cluster heads to extend the network lifetime. QPSO was applied to select the optimum cooperative coalitions of each hop in the routing path.
With consideration of mobility
The preceding studies overlooked node mobility, a critical aspect in multi-hop wireless networks where devices and nodes dynamically moved in a network environment. Mobility plays an important role in various wireless communication scenarios, notably in 6G cellular networks, ad-hoc networks, vehicular networks, and UAV networks. Several works have employed QC algorithms to address the challenge. For instance, the optimized link state routing protocol, designed for mobile ad-hoc networks, was explored in [170] and [171] from the same researchers, both leveraging QGA approaches. The dynamic topology inherent in mobile ad-hoc networks pose significant challenges in finding and sustaining an optimal end-to-end path. However, the heuristic Q-learning strategy in RL enabled dynamic adjustment of routing paths through interactions with the surrounding environment. An enhanced QEA strategy has incorporated a novel Q-learning strategy, optimizing the selection of multi-point relays [170]. This improvement addressed the limitations of traditional protocols and demonstrates properties of convergence and global optimization. Other network aspects were additionally considered in [171], such as the delivery rate of data packets, and overhead of network topology control.
Additional networks, including UAV networks, were examined with a focus on node mobility in [172] and [173], utilizing QPSO and quantum ant lion algorithm, respectively. Specifically, [172] combined the differential evolution algorithm with QPSO to further enhance the performance of both algorithms. The primary objective was to describe a general route planning strategy. The route planning for UAV on the sea was formulated as an optimization problem. A route planner for UAV was designed to generate a safe and flyable path in the presence of different threat environments. [173] presented an approach for designing an energy-efficient neuro-fuzzy cluster-based topology construction with a metaheuristic route planning algorithm for UAVs. The model utilized three input parameters: residual energy in UAVs, average distance to nearby UAVs, and UAV degree, for selecting cluster heads and constructing the topology. Additionally, quantum ant lion based metaheuristic route planning was employed to select an optimal set of routes for inter-cluster UAV communication. However, the proposal divided the problem into two independent sub-problems, potentially leading to sub-optimal solutions.
IV-B3 Applications of AI-driven QC for Routing
In recent years, AI techniques, such as RL and NNs, can be integrated with QC to enhance decision-making. AI-driven QC has emerged as a promising paradigm for addressing routing challenges to offer solutions in complex and dynamic wireless environments.
Without consideration of mobility
A framework [174] was introduced aiming at leveraging QRL to enhance the design of next-generation wireless systems. The framework, termed QRL for relay and transmit power selection, enabled sensor nodes to make optimal decisions regarding relay and transmit power selection by leveraging both present and past local state knowledge. Initially, the relay and transmit power selection problem was formulated as MDP. Subsequently, the QRL optimization aspect of the MDP was addressed, focusing on jointly optimizing energy consumption and throughput to maximize network utility. To solve this MDP problem, a QRL algorithm was developed, drawing upon Grover’s iteration. The algorithm enabled efficient decision-making within the network, facilitating enhanced performance and resource utilization. Especially, the simulation was implemented using quantum virtual machine hosted on Rigetti’s Forest platform, which provided an additional approach to validating QC algorithms. Quantum Multiple Agent Reinforcement Learning (QMARL) was also applied in tackling routing problems. The work [175] introduced a quantum-based solution for MARL tailored for autonomous control and coordination of multiple robots within smart factory environments. The delivery problem shared similarities with the routing problem. It presented an innovative QMARL framework, featuring centralized training and decentralized execution. The framework leveraged parameter sharing on policy, design of variational quantum circuits, and 2-variables dense encoding. In particular, the proposed system enabled the QDRL server to make distributed and sequential decisions for each autonomous mobile robot. These decisions revolved around determining the optimal allocation of goods, including the number of goods to carry in each autonomous mobile robot and when to request quality control. The objective was to prevent both overflow and underflow scenarios during the delivery of goods by each autonomous mobile robot.
With consideration of mobility
The previous studies focusing on AI-driven QC overlooked the consideration of node mobility in wireless networks. This aspect was addressed by a limited number of studies, such as [176]. Multiple mobile UAV platforms acted as mobile BSs to gather data from other devices. Note that trajectory planning of multiple UAVs fell within the domain of routing, as data from devices were relayed by multiple UAVs to reach the destination. The proposed QMARL-based algorithm offered a notable advantage in reducing the action dimensionality. By employing project value measures, it effectively compressed the size into a logarithmic scale. This reduction proved beneficial for enhancing the convergence of training in large MARL scenarios.
IV-C Task Offloading
IV-C1 Overview
Research in edge computing task offloading encompasses a spectrum of crucial topics [179, 180]. Efficient task partitioning and scheduling, ensuring optimal allocation of tasks between edge devices and the cloud, form a foundational aspect. The effective management of limited resources on edge devices, spanning computing, storage, and communication, is a critical consideration [181].
TABLE X: Summary of QC Works on Task Offloading in Wireless Networks
QC category Network feature Work Network structure Highlights Optimization objectives QC alg. Control manner
Conventional QC Without Consideration of Mobility [182] The network offers $K$ types of network function virtualization service. Virtual machines are divided into groups serving network function virtualizations. Formulate scheduling network function virtualization as ILP, and transform it into QUBO that is solved by QA Minimize network delay QA Centralized
[183] Multiple virtual network function instances deployed into multiple physical hosts Provide a detailed implementation, discuss the process of embedding ILP into QUBO Minimize overall resource utilization QA Centralized
[184] A computing graph network with vertices as computing servers, and edges as vertex connection links Leverage linearization and Benders’ decomposition to convert mixed ILP problem into an ILP master problem and a linear programming subproblem, present a hybrid Quantum Benders’ decomposition algorithm Minimize total learning costs of CPU iterations QA Centralized
[185] IoV: RSUs, edge servers close to RSUs, and a cloud server. RSUs perceive environment. Edge servers and the cloud process data. Introduce a quantified strategy to tackle the edge server placement problem. Employ Pareto QGA to obtain appropriate solutions Maximize the ratio of the amount of data sent to edge servers, minimize load variance of edge servers and task waiting time QGA Centralized
[186] Three network layers: sensor layer, edge layer and cloud layer Introduce task offloading to improve task execution speed for irrigation and fertilization Minimize total task completion time in a long term QGA Centralized
[187] Three layers: devices, edge nodes, and the cloud Utilize a double-hashing technique for decoding process of quantum particles Maximize throughput, minimize delay and energy consumption, and balance system load QPSO Centralized
[188] Three layers: devices, edge nodes, and the cloud Map task scheduling strategies to positions in the search space. The value in each position denotes computing resource. Maximize user satisfaction QPSO Centralized
[189] Three layers: devices, edge nodes, and the cloud Propose an improved QPSO to solve the optimization problem for optimal task offloading Minimize energy consumption QPSO Centralized
[190] IoV: offload tasks from vehicles to clouds, edge nodes, or other vehicles Establish a multi-task offloading model based on dynamic edge–end collaboration Minimize task offloading delay QPSO Centralized
[191] Two layers: devices (tasks), edge nodes (agents) Accomplish task offloading and scheduling to agents aimed at finding the optimal coalition Maximize coalition stability and user satisfaction QPSO Centralized
With consideration of Mobility [192] IoV: R2I and V2V communications Maintain population diversity; Introduce crossover operator to exchange information among individuals Minimize delay cost and energy consumption QPSO Centralized
[193] Three-tier MEC based heterogeneous 5G networks with macro, pico and femto Propose an uplink resource allocation with RF energy harvesting in a mobility scenario Maximize network throughput QPSO Centralized
AI-driven QC Without consideration of Mobility [194] WiFi/5G inter-working vehicle network: RSUs, one BS, and vehicles Design a QRL method to obtain real-time optimal resource management policy Minimize system latency QRL Centralized
[195] M IoT devices connecting to a cloud server equipped with CPUs, GPUs and QPUs Design a dynamic computation offloading strategy in which both classical and QC resources coexist Minimizes total task-offloading latency QDRL Centralized
[196] An edge computing system: mobile users and quantum edge servers Formulate a classical and quantum computation offloading problem as a mixed ILP, solved by hybrid continuous-discrete MDRL Minimize total offloading and execution cost Quantum multiple-agent DRL Distributed
[197] MEC networks: multiple IoT devices and edge servers Employ quantum state representations to transform MDP into a QC problem Maximize system energy efficiency QDRL Centralized
[198] MEC networks: multiple IoT devices and edge servers Combine QC and ML to achieve exploration and exploitation trade-off via quantum parallelism Maximize system energy efficiency while guaranteeing transmission latency requirements QDRL Centralized
[199] VRs have computing capacity to perform immersive viewpoint rendering, associated with an edge AP to access VR servers Integrate wireless VR QoE and content correlation together with computation offloading Minimize the total energy consumption QRL Centralized
[200] VRs associated with multiple edge AP to receive videos and feedback from the server Integrate content correlation, block consensus, fluctuating channel conditions, QoE, and the blockchain to offload tasks for security Minimize energy consumption QDRL Centralized
[201] UAVs offloading tasks to BS, with limited resources, to construct Metaverse Propose a QDRL algorithm to render UAV images to construct Metaverse Maximize the benefits for Metaverse-service provider QDRL Centralized
TABLE XI: Details of QC Algorithms in Task Offloading
Work Alg. Orth. QC Quantum computer Platform No. qubits Mobility
[182] QA D-wave Unknown
[183] QA D-wave Unknown
[184] QA D-wave Unknown
[185] QGA Computer simulation 0
[186] QGA Computer simulation 0
[187] QPSO Computer simulation 0
[188] QPSO Computer simulation 0
[189] QPSO Computer simulation 0
[190] QPSO Computer simulation 0
[191] QPSO Computer simulation 0
[192] QPSO Computer simulation 0
[193] QPSO Computer simulation 0
[194] QRL Computer simulation 0
[195] QDRL Computer simulation 0
[196] QMADRL Unknown [1000, 5000] per edge server
[197] QDRL Computer simulation 0
[198] QDRL Variational quantum circuits Unknown
[199] QRL Computer simulation 0
[200] QDRL Computer simulation 0
[201] QDRL Computer simulation 0
IV-C2 Applications of Conventional QC for Task Offloading
Without consideration of mobility
Traditional task offloading, which involves scheduling function virtualization services to users, was explored in [182, 183]. Specifically, in [182], the network provided $K$ types of network function virtualization services, with virtual machines organized into groups dedicated to serving specific network functions. The scheduling problem for network function virtualization was initially framed as an Integer Linear Programming (ILP) challenge, subsequently converted into a QUBO model, and addressed by QA. Similarly, [183] addressed the task of offloading virtual network functions within a simple network setup. It provided a detailed implementation introduction, discussing the process of embedding the ILP problem into a QUBO format. It included the formation of the Q-matrix, crucial for solvability by QA, and the mapping of quantum hardware to the problem domain.
With the evolution of 5G and beyond, edge computing emerges as a compelling paradigm aimed at delivering services to users by offloading computation tasks from users and cloud infrastructure to edge nodes situated in close proximity to users. In response to the challenge posed by limited edge server placement in IoV networks, a method called quantum edge server placement is developed in [185]. The method employed dual encoding, utilizing both binary and quantum encoding techniques, to represent the locations of edge servers. Leveraging the non-dominated sorting genetic algorithm, it effectively explored potential placement solutions. By employing an assessment function, it derived optimal placement schemes, thereby improving the efficiency of edge server deployment in IoV scenarios. However, it is worth noting that it did not account for network dynamics, such as the impact of vehicle mobility. QGA-based task offloading also found applications in intelligent agriculture, including irrigation and fertilization tasks [186]. Quantum operators utilized rotation gates and non-gates to update the matrix, mitigating the algorithm complexity by eliminating query operations from the rotation angle table during quantum rotation gate updates. However, it is worth noting that in practice, the latency requirements for drip irrigation and fertilization tasks were not stringent, and the demand for computational resources in these applications tended to be low. Consequently, these factors resulted in less attention being given to finding optimized solutions for such scenarios.
Another conventional technique for task offloading is QPSO, widely applied in service placement within IoT networks [187], as well as in multi-task applications [189, 190, 191]. A QPSO method is proposed specifically for IoT task offloading in edge computing systems in [187]. The problem was formulated mathematically, with quantum particles efficiently designed to handle the entire IoT task offloading process. These particles guaranteed the generation of complete and valid solutions to the problem. Notably, the decoding process of quantum particles employed a novel double-hashing technique, ensuring uniform task offloading to the edge nodes. Furthermore, the fitness function was designed considering four distinct objectives: maximizing throughput, minimizing delay, reducing energy consumption, and balancing system load.
Federated learning, a method of distributing computational tasks across multiple decentralized servers [184], can be considered a variant of task offloading. The approach addressed the challenge of training machine learning models across a distributed network without centralized data aggregation. The process involved formulating a joint participant selection and learning scheduling problem as a mixed ILP problem, aimed at minimizing the total learning costs. To tackle this optimization problem, a hybrid quantum-classical Benders’ Decomposition algorithm was proposed. By combining classical and QC methodologies, the approach aimed to efficiently solve the optimization problem associated with multi-modal federated learning in distributed networks.
The previous studies fail to account for the impact of multiple tasks on system performance. For instance, in [187], while mathematical models involved multiple users with multiple tasks, there was no mutual influence among them. The task offloading problem was mathematically formulated as a mixed-integer nonlinear programming (MINLP) model in [188]. The formulation took into account constraints such as deadlines and resource capacities, with the objective of maximizing user satisfaction. Within the proposed QPSO algorithm, task scheduling solutions were mapped to positions in the search space using real number coding. Here, the value in each position denoted the computing resource to which the corresponding task is assigned. The algorithm adopted a population evolutionary approach of QPSO to iteratively search for the global best solution.
The study presented in [189] introduced an innovative solution to the task offloading within MEC environments characterized by users and servers. In this study, the task offloading strategy was reformulated as an energy optimization problem. To address this challenge, an enhanced QPSO algorithm was proposed, aimed at efficiently solving the optimization problem for optimal task offloading. The effectiveness of the proposed QPSO-driven offloading strategy was demonstrated through its convergence. A bound state described by the probability density function, appeared at any interval of the completely searchable space with a certain probability. However, it was noteworthy that the impact of different tasks remained relatively weak, as multiple tasks from various users competed for the limited resources available at edge nodes. Additionally, the study lacked analysis regarding the impact of mobility within MEC environments. Similar limitations were observed in the study conducted by [190] within the IoV environment. Additionally, the assumption that the next hop was predetermined in scenarios where tasks were offloaded from vehicles to other vehicles lacks practicality. The first limitation was addressed in [191]. It identified two interrelated impacts among tasks: firstly, when a complex task arrived, it was offloaded to a coalition of edge nodes for collaborative resolution; secondly, the complex task was divided into multiple sub-tasks, each offloaded to distinct coalitions. The proposal calculated the Shapley value of the optimal solution by employing a similarity metric based on the task rewards of each combined edge group.
With consideration of mobility
The previous works in this subsection overlook the mobility of end-uses or edge nodes even though in MEC or IoV scenarios. Given the rapid advancements in mobile communication technology, a diverse array of intelligent computation systems have surfaced, demanding more stringent QoS performance.
The channel resource allocation was presented in [193] in the background of task offloading in a three-tier, such as macro, pico and femto, MEC based heterogeneous 5G networks. The user equipment offloaded computation tasks to the local server, which significantly reduced the transmission latency. In the process of data transmission, multiple nodes competed for the channel resource. However, a random mobility model was designed to describe the movement of the user equipment. One approach in [192] introduced a joint offloading strategy based on QPSO for MEC-enabled vehicular networks. The strategy focused on minimizing delay costs and energy consumption by offloading tasks from vehicles to available service nodes, including service vehicles and nearby RSUs. To facilitate task offloading via V2V communication, a vehicle selection algorithm was introduced to determine an optimal offloading decision sequence. Moreover, to maintain population diversity in the QPSO algorithm, a crossover operator was employed to exchange information among individuals. However, it was assumed in the mobility model that vehicle speeds and coordinates were given in advance, presenting a practical limitation.
IV-C3 Applications of AI-driven QC for Task Offloading
The emergence and rapid development of AI have transformed QC techniques for task offloading into a higher level of smartness and efficiency through many scenarios, e.g., IoV [194] and smart healthcare [195]. Specifically, in [194], a resource management strategy was introduced to achieve the balance between between communication and computation requirements by continuously monitoring the status of local vehicles, MEC-assisted WiFi links, and cloud-assisted 5G cellular networks. It empowered vehicles to make adaptive and flexible decisions regarding task offloading and transmission. To tackle the optimization problem in the face of time-varying channel conditions, it was reformulated as a MDP, solved by QRL to derive the real-time optimal resource management policy. The approach posited a virtual quantum server for performing optimization, whereas works [195, 196] explicitly introduced quantum servers.
In [195], a dynamic computation offloading schedule was designed to minimize total latency within a hybrid setup, accommodating both classical (CPUs/GPUs) and quantum (QPUs) computing resources in cloud servers. This strategy ensured sustainable computation offloading while meeting individual latency constraints and the requisite success ratio for each task. Initially, the latency minimization problem was reformulated as a step-wise mixed-integer non-convex optimization task utilizing Lyapunov techniques. Subsequently, a DQN was employed for computation offloading mode selection, CPU/GPU mode or QPU mode, for cloud servers, addressing the optimization problem. Additionally, quantum edge servers were established in [196]. It formulated a classical and quantum computation offloading problem as a mixed-integer programming model, which was tackled using a multi-agent hybrid continuous-discrete DRL algorithm.
QC coupled with DRL techniques also found extensive applications in addressing task offloading challenges across various scenarios, including IoT [197, 198], VR [199, 200], and Metaverse environments [201]. In [197], a MEC-enabled IoT framework was proposed, integrating QC capabilities. The architecture enabled IoT users to make sequential offloading and transmission decisions in an adaptive and flexible manner. The study formulated a joint optimal stochastic computation task offloading and dynamic resource allocation problem aimed at maximizing system energy efficiency. The optimization considered factors such as transmission latency, energy consumption, time-varying channel conditions, and data buffer constraints. Quantum state representations were employed to transform the continuous-time MDP problem into a QC optimization, which was effectively addressed by the proposed QDRL algorithm. The research team also extended their solution into 6G MEC based IoT system [198], but applying variational quantum circuits.
Representative applications in 6G, VR and Metaverse, presented challenges in transmitting resources for large-volume streaming data, excessive computing resources for encoding and decoding operations, and extreme energy consumption in transmitting and receiving devices. QC coupled with DRL techniques emerged as a promising approach to address these challenges [199, 200, 201]. The study [199] presented an energy-aware resource management strategy for VR-supported industrial IoT. It addressed rendering task offloading and resource allocation challenges considering factors like content correlation, fluctuating channel conditions, and VR QoE. To handle this problem effectively, it was transformed into a sequential decision-making scenario under varying channel conditions over time. By defining a reward function that encapsulated energy consumption and VR QoE metrics, the optimization problem was formulated as a MDP, which was then tackled using the QRL algorithm to maximize the long-term system reward. The researchers extended their work to enhance resilience against malicious attacks by integrating the blockchain technique into the system, particularly in the context of VR-enabled medical treatment applications in 6G [200]. Additionally, they applied the QDRL algorithm instead of the QRL algorithm to enhance optimization efficiency. Another innovative application of 6G, Metaverse, leveraged DRL-driven QC algorithms to facilitate seamless immersion between reality and virtuality. Service subscribers accessed a real-time virtual world, constructed from images captured by on-site UAVs, within the Metaverse framework [201]. A key challenge arose in determining the optimal BSs to which UAVs offloaded their tasks to maximize benefits, particularly when BSs have limited bandwidth and computation resources. Specifically, to tackle the challenge posed by new environments or re-deployments, where traditional RL methods often encountered model collapse, collective learning was introduced. This approach enabled UAVs to leverage shared intelligence, facilitating rapid adaptation to new environments. The work can be improved by considering UAV mobility to achieve a more efficient system.
In conclusion, QC offers valuable insights into task offloading, particularly in the context of evolving networking technologies like 6G, IoV, and UAV networks. Detailed summaries of these works are presented in Table X, with further elaboration on their QC algorithms available in Table XI. It is worth noting that task offloading within these scenarios typically involves distributed control and node mobility. However, only a few works provided solutions addressing these aspects, showed in Table XI. The primary reason for the first characteristic is current high overhead of deploying quantum machines. To the best of our knowledge, mobility remains an open challenge for task offloading integrated with QC.
IV-D Content Caching
IV-D1 Overview
In the fast-paced world of digital content consumption, content caching, a technique widely employed by content delivery networks (CDNs), web servers, information centric network, etc., plays a pivotal role in optimizing the delivery of online content, profoundly influencing the overall user experience [202, 203]. It involves the strategic storage of frequently accessed content or computation results on servers distributed geographically, ensuring that users can access information with minimal latency, cost or other preferred objective. This practice significantly increases responsiveness and network bandwidth utilization.
Content caching scenarios are prevalent across various applications, spanning from edge computing [204], 5G cellular networks [205, 206], IoT [207] to video streaming [208, 209]. With the development of QC algorithms, the solutions leveraged QGA [204], QDL [210], and QRL [209] to tackle problems. The inherent technical challenges in content caching stem from several key factors. First, ensuring effective caching cooperation among different network entities is crucial for optimizing the distribution of cached content and maximizing overall network performance. Second, resource limitations, such as limited storage capacity and bandwidth availability, pose significant challenges in efficiently managing cached content and satisfying user demands. Third, the dynamic nature of network traffic and user preferences requires adaptive caching strategies that can adjust in real-time to changing conditions. Fourth, the heterogeneity of network devices and their capabilities introduces complexities in coordinating caching decisions across different devices with varying computational and storage resources. Finally, the distributed control manner of content caching systems necessitates robust algorithms and protocols for coordinating caching activities and ensuring seamless integration within the network architecture.
TABLE XII: Summary of QC Works on Content Caching in Wireless Networks
QC category Work Network structure Highlights Optimization objectives QC alg. Control manner
Conventional QC [204] 5G cellular networks: a macro, picos and users Introduce an adaptive weighted cost function to combine delay and energy cost Minimize delay and energy cost QGA Centralized
[205] 5G: a macro BS, multiple small-cell BSs that can be activated as either a pico or a femto Design a joint caching and activation policy to improve energy efficiency Maximize system energy savings QEA Centralized
[206] 5G: a core gateway, multiple eNodeB and users Propose a hierarchical cooperative caching policy among a core gateway and eNodeBs to realise energy efficiency Minimize total transmission energy consumption QEA Centralized
[208] 5G: Cache and relay bit-rate video streaming with assistance of UAVs Propose a QEA algorithm to optimize joint caching and user association for adaptive video streaming Minimize total content delivery delay QEA Centralized
AI-driven QC [210] Fog-computing assisted networks: a centralized cloud, fogs and users Integrated DL into QC to address video caching with a detail process of training a QDL model Update QDL model parameters according to the dataset QDL Centralized
[207] Edge computing assisted networks: a BS, edge servers and IoT devices Optimize resource allocation, content caching, and task offloading in dynamic large-scale IoT networks Maximize energy efficiency QRL Centralized
[209] Edge computing assisted networks: a BS, edge servers and users Optimize content caching and spectrum allocation, considering dynamic content popularity, transmission method, and requirements on QoE Minimize content delivery delay QRL Centralized
[211] Edge computing assisted networks: a BS, edge servers and users Integrate DRL into QC to address task offloading and content caching. Maximize the caching benefit QDRL Centralized
IV-D2 Applications of Conventional QC for Content Caching
To alleviate backhaul link congestion and improve resource utilization in 5G ultra-dense cellular networks consisting of a macro, multiple picos and users, it is a traditional issue to cache content cooperatively in picos with resource limitations as illustrated in [204]. This study focused on an energy–delay tradeoff problem in collaborative edge computing-assisted and energy harvesting-powered ultra-dense networks. It regarded delay and energy as two types of cost and introduced a weighted cost function to transform the tradeoff problem into a cost minimization problem. An alternating optimization based on an improved QGA was proposed to solve the cost minimization problem. In the alternating optimization, the proposal divided the cost minimization problem into two sub-problems, adaptive tuning weight sub-problem and caching decision sub-problem. It improved QGA in terms of repairing unfeasible solutions and adaptively updating quantum genes. The problem arose from the fact the solutions derived from the individual sub-problems can not ensure the optimality of the original problem.
One of important features in 5G cellular networks is heterogeneity, i.e., small-cell base stations can be activated when necessary to be either a pico-cell base station or femto-cell base station [205], and cooperative caching performs in core gateways and BSs [206]. A joint caching and activation mechanism in [205] was designed to maximize system energy savings, while a cooperative caching policy was proposed in [206] to reduce the duplicate transmission and improve the green content delivery. Both two methods formulated the issue as an ILP problem, which were solved by QEA. Another heterogeneity came from assistance of UAVs for edge caching in 5G cellular networks, which had also received extensive attention from both industry and academia [208]. UAV-assisted edge caching is a promising technique to improve content transmission and enhance delivery efficiency by caching video contents closer to users. The study in [208] adaptively adjusted video quality based on time-varying network conditions and different user preferences according to proposed QEA. However, the work can be further enhanced by leveraging UAV mobility to dynamically adapt to user mobility patterns.
IV-D3 Applications of AI-driven QC for Content Caching
AI-driven QC can significantly enhance the efficiency of content caching by leveraging its advantages in computing [210, 207, 209, 211]. In [210], DL was integrated into QC to address the traditional challenge of video caching with a detail process of training a QDL model. Specifically, a DL agent utilized the self-organizing maps algorithm to prioritize caching contents. The prioritized contents were stored in a quantum memory module using a two-level spin quantum phenomenon. Following the selection of the most suitable lattice map via self-organizing maps, data points below a specified threshold were mapped onto the data frame to retrieve the videos. These videos were considered high-priority for training based on the input data provided in the dataset. Similarly, the models for the other level prioritized content can be trained accordingly.
Embedding RL into QC [207, 209] was another effective approach to solving complex problems in content caching, especially for the problems in sequential decision processes. In [207], there were two sub-problems of task offloading and content caching with limited resources in a traditional edge computing assisted network. The network architecture comprised a BS, edge servers, and IoT devices. The proposed approach aimed to optimize dynamic resource allocation, content caching strategy, and computation offloading policy to enhance energy efficiency in dynamic large-scale IoT networks. A QRL based algorithm was introduced to address the challenges posed by complex and dynamic environments, varying content popularity, and user preferences. However, there were two shortcomings in the proposal. First, the rationale behind jointly addressing the offloading and caching problem was unclear, lacking practical justification. Secondly, the solution relied on integer-based decisions for offloading and caching, potentially limiting its flexibility and scalability. QRL was also applied in the application of caching video streaming to improve QoE for users [209]. Its proposal jointly optimized content caching and spectrum allocation to minimize content delivery delay, taking into account time-varying content popularity, transmission method selection, and different requirements on EoE in a similar wireless network structure with [207]. The second shortcoming in [207] also arose in [209] from the binary caching and offloading decisions in its proposal. To address the same problem as [207], the study [211] proposed a solution by integrating DQN into QC, which involved the joint optimization of task offloading and content caching. However, a limitation arose as the two sub-problems were optimized independently, potentially leading to sub-optimal solutions for the original problem.
In conclusion, QC algorithms are widely applied in the area of content caching, as presented in Table XII. There are still several issues in existing literature. Many related works implement their QC algorithms in computer simulations duo to the current developing stage of quantum computers. Additionally, current solutions often employ centralized control methods, primarily because computation services typically take the form of quantum clouds. Lastly, similar to the challenge in task offloading, node mobility remains an open issue when considering content caching.
V Miscellaneous Issues
The first section explores QC impact on wireless security, focusing on encryption, authentication, and privacy techniques like federated learning and blockchain. The second part shows the QC and other emerging methods for pinpointing wireless device locations. The third section discusses techniques for optimizing video content delivery over wireless networks, addressing challenges such as the optimization to improve streaming performance.
V-A Security and Privacy
On the one hand, quantum communication stands out as a promising frontier in securing sensitive information against eavesdropping and unauthorized access. Unlike classical cryptographic techniques, quantum communication leverages the principles of quantum mechanics to ensure the security of transmitted data. For example, quantum communication protocols like quantum key distribution [212], exploit the inherent properties of quantum particles, such as superposition and entanglement, to establish secure communication channels. On the other hand, QC enables quantum computers to perform complex computations at exponentially faster rates than classical computers, making them well-suited for cryptographic applications, such as secure coding such as secure URLLC in 5G and beyond cellular networks [213, 214].
Attacks on wireless pilot signals pose a significant threat to URLLC services. In grant-free URLLC scenarios, accurate allocation of distinct pilot resources to multiple users sharing the same time-frequency resource is crucial for the next-generation Node B to accurately identify these users under access collisions. This precise allocation is essential for maintaining accurate channel estimation, which is necessary for ensuring reliable data transmission. However, this procedure is susceptible to attacks targeting pilots. The research in [213] introduced an approach utilizing quantum learning to implement nonrandom superimposed coding for encoding and decoding pilots across multidimensional resources. The method enabled rapid learning and precise elimination of attack uncertainties. Specifically, pilots for uplink access by multiple users were encoded as distinct sub-carrier activation patterns. The Node B then decoded the desired pilots from observed sub-carrier activation patterns, which represented a superposition of sub-carrier activation patterns from accessing users. This was achieved through a joint design involving attack mode detection and user activity detection via a quantum learning network. The study in [214] solved the same security problem as [213] but in the area of random structure coding. It introduced an approach to encoding and decoding pilot signals using a random structured code across multidimensional physical resources. The proposed code employed random encoding with minimal structure to mitigate the impact of attacks. By leveraging group spatial channel features, the decoding process was modeled as a computational trap, leading to a security challenge akin to random computing with redundancy. To address this, the study utilized a quantum algorithm to learn the computational trap model, enabling the removal of computational redundancy and elimination of dispersed attacks. The proposal demonstrated the existence of a quantum black-box model corresponding to the computational trap and derived a precise expression for computational performance.
QC can also integrate with existing popular security and privacy techniques like blockchain [215] and federated learning [216, 217, 218] to enhance network performance. Inspired by classical federated learning techniques, the work in [216] aimed to enhance data privacy by aggregating updates from local computations to share model parameters. In pursuit of approximate optima in the parameter landscape, it introduced an extension of the traditional variational quantum algorithm. The variational quantum algorithms were also applied in future 6G space-air-ground integrated networks to address the complexity emerging from vast datasets and ML models aiming at data security and privacy [218].
V-B Localization and Tracking
Existing localization and tracking solutions often require continuous communication and processing, leading to significant energy consumption, particularly in battery-constrained devices [219]. Additionally, it faces challenges in scaling to support a large number of devices simultaneously, leading to increased computational complexity and resource consumption. QC facilitates simultaneous processing of multiple possibilities, which accelerates and enhances localization and tracking tasks. This capability finds applications in various scenarios, including tracking mobile user trajectories [220], managing wireless sensor networks [221, 222], and coordinating multiple UAV networks [223]. User trajectory prediction in [220] involved two components: reservoir computing and QC. Compared to training simple recurrent neural networks, reservoir computing was computationally more efficient because only the output layer weights were trainable. However, achieving optimal performance in reservoir computing requires careful selection of reservoir weights to create complex and nonlinear dynamics. QC was employed to generate such complex reservoir dynamics, mapping input time series into a higher-dimensional computational space consisting of dynamical states. Once high-dimensional dynamical states were obtained, a simple linear regression was applied to train the output weights, enabling efficient prediction of mobile users’ trajectories.
QC also finds extensive applications in localization and tracking within sensor networks, particularly addressing scalability challenges. The approach in [221] enhanced the differential crossover in QPSO to address nonlinear localization problems influenced by node density and coverage in IoT. By utilizing the differential evolution operator, it mitigated group movements within narrow ranges, thus preventing convergence to local optima and enhancing global search capabilities. The cross operator facilitated information exchange among individuals within a group, while exceptional genes were moderately retained to sustain the evolutionary process, thereby incorporating both proactive and reactive features. In contrast to traditional location scenarios as presented in [221], [222] addressed the localization with multiple mobile robots equipped with sensors aiming to locate a specific source node. By modeling each robot as a double integrator, [222] converted the source location challenge into a multi-robot path planning task, considering constraints such as collisions between robots and obstacles, as well as collisions among robots. A framework for the mobile policy of the robots was devised. And an adaptive weight strategy, a swarm timely update mechanism, and a leading-following behavior were integrated into QPSO. These enhancements aimed to enable multiple robots to adhere to the mobile policy effectively, ensuring robustness and efficiency in their search.
QC, paired with AI, presents new ways to tackle location and tracking tasks in UAV networks [223]. In [223], a DRL approach was proposed for controlling vertical take-off and landing aircraft in the presence of wind disturbances. The tracking control problem was framed as MDP. To enhance control accuracy despite wind disturbances, various wind fields were incorporated into the learning environment to broaden exploration and mimic the impact of winds on flight control. To ensure both tracking accuracy and practical implementation, a quantum-inspired experience replay strategy was introduced based on QC principles. This strategy featured a preparation operation scheme to foster exploration and accelerate convergence, along with a depreciation operation method to enhance sample diversity and bolster system robustness.
Previous efforts to locate and track devices relied on sensors or GPS, which was practical for scenarios like indoor person tracking with only a cellphone. Existing solutions for such an application typically utilize received signal strength (RSS) to analyze the location, but faces a challenge of scalability to worldwide and algorithm complexity. The work [224] provided a fast matching quantum fingerprinting algorithm by calculating the quantum cosine similarity between online RSS vector at an unknown location and each RSS vector in the cellular fingerprint. Its efficiency was validated in the testbed IBM Quantum Machine Simulator and cellular BSs. While it is an innovative approach, the matching accuracy can be further enhanced by utilizing channel state information (CSI) instead of RSS. However, acquiring CSI presents a practical challenge due to hardware limitations. Moreover, RSS, while commonly used, is susceptible to noise factors like pedestrian interference, which is not accounted for in the solution.
V-C Video Streaming
Video serves as a primary means of conveying information through visual and auditory media and finds widespread applications in entertainment, education, and information dissemination. A typical video system includes several components such as video encoding/decoding [225], network transmission [226], and buffer [227], etc. differing significantly from conventional data transmission. Additionally, emerging technologies such as multi-view videos [228], free-viewpoint videos [229], VR/360 videos [230], and point cloud videos [231, 232] have surfaced in recent years. These innovations allow users to choose the viewing angle, enabling them to selectively watch specific parts of the video during the viewing process, for example, in the case of 360 videos/point cloud videos, where users can switch perspectives over time.
In wireless communications and networking, one classical issue is adaptive bitrate (ABR) control in order to ensure the video streaming quality. The work [233] maximized the user QoE by increasing the average bitrate and decreasing the video rebuffering events. Specifically, the proposal formulated the control problem into a QUBO model from three aspects: video quality, bitrate change, and buffer condition, which was solved in the platform Digital Annealer. Representative applications, VR and Metaverse in 6G and beyond, also employed QC algorithms to improve performance. Ultrahigh viewport rendering demands and excessive terminal energy consumption restrict the application of wireless VR. The work [234] considered content correlation between VR equipments, fluctuating channel conditions, and VR QoE, by formulating the joint problem of the viewport rendering offloading, computing, and spectrum resource allocation. The formulation was solved by proposed RL algorithm and quantum parallelism was integrated into the RL to overcome the problem of high data dimension.
VI Challenges and Open Issues
VI-A QC in Wireless Communications and Networking with Multiple Variables
QC has been applied in wireless communications and networking, primarily focusing on individual variables. However, tackling optimization problems involving multiple variables poses a more formidable challenge. Addressing multiple variables in wireless systems presents significant hurdles for optimization and control. The exponential increase in dimensionality of optimization problems renders them computationally intensive and impractical for real-time decision-making. Balancing conflicting objectives, such as maximizing throughput while minimizing latency and interference, becomes intricate. Moreover, dynamic environments exacerbate the complexity, introducing uncertainty in predicting optimal configurations. Complex interactions among variables, compounded by resource constraints, further compound the challenge. Efficient algorithms that integrate optimization theory, signal processing, and machine learning are essential to effectively tackle these challenges.
On the other hand, AI-driven quantum algorithms sometimes require datasets for training, which is currently hindered by insufficient available datasets.
VI-B QC in Wireless Communications and Networking for Particular Services
Introducing QC into wireless communications and networking for specific services poses several challenges. For instance, services like video streaming demand stringent requirements beyond conventional wireless systems, such as low latency, high throughput, and reliable transmission. Optimizing these parameters using quantum algorithms becomes more challenging due to the complexity introduced by the additional service requirements. Quantum algorithms must navigate the trade-offs between different objectives, such as maximizing video quality while minimizing latency and packet loss. Furthermore, the dynamic nature of video traffic adds an extra layer of complexity, requiring adaptive optimization strategies that can efficiently allocate resources in real-time. Additionally, ensuring the security and privacy of video data transmission becomes more critical with the integration of QC, necessitating robust encryption schemes and secure communication protocols that can withstand quantum attacks. Overall, integrating QC into wireless communications and networking or specific services like video streaming presents challenges in balancing competing objectives, adapting to dynamic traffic patterns, and ensuring robust security measures, requiring innovative solutions tailored to the unique requirements of each service.
VI-C Understanding and Utilizing QC in Communications and Networking
While QC has demonstrated remarkable efficiency and achieved excellent results in some systems, their widespread application still faces numerous challenges compared to techniques like convex optimization and DRL. Particularly, QC, as a novel approach distinct from conventional methods, presents difficulties in understanding its operational principles. Unlike traditional algorithms whose operations are more easily comprehensible, the workings of QC involve intricate quantum mechanics concepts, making them less intuitive to grasp. Bridging this comprehension gap and facilitating the integration of QC into wireless systems pose significant challenges. Educating individuals about this technology and its applications in wireless communications and networking requires simplifying complex quantum concepts into more understandable terms and providing accessible learning resources. Additionally, developing user-friendly tools and interfaces that abstract away the complexity of quantum operations can aid in democratizing the utilization of QC in wireless systems. Overcoming these challenges demands collaborative efforts from researchers, educators, and industry professionals to demystify QC and foster its practical implementation in wireless communications and networking effectively.
VI-D QC Cloud Services in Wireless Communications and Networking
As QC becomes increasingly available to the public through cloud services, there is a growing need to ensure these services can be effectively supported in wireless networks in the near future [235]. Factors such as communication latency and stability will be critical considerations in meeting network requirements. Furthermore, the integration of QC into the classical computing pipeline will involve determining which components are most suitable for quantum processing to maximize benefits. This includes addressing the overhead associated with transitioning between classical and QC, a crucial aspect for optimizing overall efficiency and performance.
VII Conclusions
The advent of QC technology signals a significant revolution in wireless communications and networking. To facilitate the exploration of its applications in wireless communications and networking, we provided a tutorial on QC, covering its basics, characteristics, and popular QC algorithms. Additionally, we introduced the applications of QC in communication and networking, followed by its applications in miscellaneous areas such as security and privacy, localization and tracking, and video streaming. We believe this is a very promising direction and have discussed the remaining open issues.